### v3

# Exploring AI-based System Design for Pixel-level Protected Health Information Detection in Medical Images

Abstract
De-identification of medical images is a critical step to ensure privacy during data sharing in research and clinical settings. The initial step in this process involves detecting Protected Health Information (PHI), which can be found in image metadata or imprinted within image pixels. Despite the importance of such systems, there has been limited evaluation of existing AI-based solutions, creating barriers to the development of reliable and robust tools. In this study, we present an AI-based pipeline for PHI detection, comprising three key components: text detection, text extraction, and text analysis. We benchmark three models—YOLOv11, EasyOCR, and GPT-4o— across different setups corresponding to these components, evaluating the performance based precision, recall, F1 score, and accuracy. All setups demonstrate excellent PHI detection, with all metrics exceeding 0.9. The combination of YOLOv11 for text localization and GPT-4o for extraction and analysis yields the best results. However, this setup incurs higher costs due to GPT-4o’s token generation. Conversely, an end-to-end pipeline that relies solely on GPT-4o shows lower performance but highlights the potential of multi-modal models for complex tasks. We recommend fine-tuning a dedicated object detection model and utilizing built-in OCR tools to achieve optimal performance and cost-effectiveness. Additionally, leveraging language models such as GPT-4o can facilitate thorough and flexible analysis of text content.
Keywords: de-identification, Protected Health Information (PHI) detection, medical imaging, Large Language Model (LLM)
Statements and Declarations
Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
Competing interest
The authors declare they were employed by Bayer AG during the time of the research. No other financial or non-financial interests related to the work submitted for publication are disclosed.
Author contributions
All authors contributed to the conception and design of the study. Tuan Truong was responsible for preparing materials, as well as for data collection and analysis. Ivo Baltrulschat and Matthias Lenga provided valuable input through discussions and contributed to the analysis of the experimental results. Mark Klemens and Grit Werner offered insights through their clinical discussions. All authors have read and approved the final manuscript.
Ethics approval
This research study was conducted retrospectively using human subject data made available in open access. Ethical approval was not required, as confirmed by the license attached to the open-access data.
Consent to participate
Not applicable
Consent to publish
Not applicable
Data availability
The datasets used for training and evaluation of the pipeline are available upon request.
Acknowledgements
The authors would like to thank the Bayer team of the AI Innovation Platform for providing computing infrastructure and technical support. Furthermore, we would like to thank Subrata Bose for fruitful discussions and valuable feedback.
1 Introduction
Protected Health Information (PHI) is health-related data that can identify an individual, often found in medical records, images, or any documents and data elements generated during participation in healthcare services. Before sharing data, e.g., for research purposes, it is crucial to de-identify or pseudonymize it to minimize the risk of re-identification. This de-identification step is mandatory for studies to comply with the Health Insurance Portability and Accountability Act (HIPAA) [1] and the General Data Protection Regulation (GDPR) [2], and it is the responsibility of the data provider. De-identification is a complex process that involves many data types and formats, typically including a detection step followed by the redaction of the identified PHI. Medical images are a rich source of PHI, which can be embedded as metadata (e.g., in DICOM headers) or burned-in text within pixel data. The automatic detection of pixel-level PHI is challenging due to the diversity in imprint content, locations, font sizes, color schemes, annotations, and machine overlays. As the overall volume of medical data increases, so does the risk of incomplete PHI redaction, potentially leading to data breaches, which imply significant negative consequences for affected individuals and the healthcare sector, as well as high costs for responsible entities [3]. The importance of reliable PHI detection at scale and automated image quality control in general is strongly emphasized by the latest medical image data platform development program launched by the Advanced Research Projects Agency for Health (ARPA-H) [4].
Previous works have employed Optical Character Recognition (OCR) technology to recognize text in medical images, leading to two primary approaches for handling burned-in text. The first approach removes all burned-in texts, regardless of content [5]. While this method effectively eliminates any potential PHI, it also discards valuable information, such as measurements and markers. The second approach applies heuristic rules to redact specific PHI imprints [6]. However, this method is tailored to particular representations of imprints in images and can struggle to generalize to new types of PHI or different representations. In addition, fine-grained classification of PHI is often preferred to coarse classification, which categorizes information simply as names or dates. For instance, not all patient identifiers or dates need redaction. Identifiers linked to the hospital system pose a high risk of identification and should indeed be redacted, while those used in clinical trials are valid for study analysis. Similarly, although all dates are classified as PHI under HIPAA guidelines, certain dates, such as those documenting patient visits, are critical for clinical trials. As the volume and diversity of medical data continue to grow, advanced automatic solutions for PHI detection are essential, particularly for managing large datasets collected retrospectively. These solutions are expected to handle edge cases, which can involve various signaling words and phrases. Moreover, research sites require flexibility to determine which entities to include or exclude as PHI based on specific study setup. Therefore, a robust PHI detection tool must incorporate mechanisms that enable easy specification and adjustment of rules to meet the needs of different studies.
Our study focuses on AI-based approaches for automatically detecting PHI as burned-in pixels. Currently, there are a few solutions that utilize vision and language models from MD.ai [7], Google [8], John Snow Labs [9], and Glendor [10]. These tools commonly integrate two key components: a vision component that detects and converts burned-in texts into machine-encoded strings, and a language component that identifies any PHI content in the extracted texts [11]. To achieve this, OCR models are utilized for text recognition, while name-entity recognition (NER) models are employed to classify the PHI content. Recently, large language models (LLMs) have become the preferred candidates for the language component due to their flexible configurability and ability to understand complex language context in NER tasks, which has proven helpful in detecting PHI in biomedical text [11]. Despite some general remarks concerning the deployment of these tools [12, 13], particularly on the non-deterministic behavior and cost of LLMs, there is a lack of comprehensive studies dissecting the setup of vision and language models within a PHI detection pipeline. In addition, as multimodal LLMs with vision capabilities are being released today [14], the question remains whether they can be used to recognize PHI through end-to-end processing of images. To address the above concerns, we make the following contributions in our study:
  • We propose an end-to-end PHI detection system comprising three modules: text localization to identify text in images, text extraction to convert pixels into machine-readable text, and text analysis to detect PHI content using advanced language models (Figure 1).
  • By leveraging publicly available medical imaging data and the HIPAA guidelines, we create two evaluation datasets comprising 1000 and 200 images, respectively. These datasets encompass various radiology modalities, including X-Ray, CT, PET, MRI, bone scintigraphy, mammography, and ultrasound. The images are synthesized with pixel-level PHI to support development and benchmarking within the research community.
  • We combine three models (i.e., YOLOv11, EasyOCR, and GPT-4o) corresponding to three previously mentioned modules in four different setups of a PHI detection pipeline (Figure 2) and compare their performance in an extensive quantitative study. In each setup, a model can assume one or multiple roles. For example, EasyOCR can perform not only text extraction but also text localization, and GPT-4o can perform end-to-end PHI detection instead of only analyzing the semantic content of extracted texts.
  • Based on the quantitative results, we provide insights into the optimal pipeline setup, considering performance, latency, and inference cost in terms of the number of generated tokens. Specifically, we explore the full capabilities of GPT-4o for both vision and language tasks, and analyze the trade-offs involved in utilizing it partially or entirely in the PHI detection pipeline.
Figure 1: Pixel-level PHI detection pipeline. The pipeline consists of three components: text localization, text extraction, and text analysis. First, the bounding box coordinates of imprints are detected. Then, the text extraction module transforms the pixel-level imprints into machine-encoded texts. In the last step, the text analysis module identifies the PHI content.
Figure 2: Proposed experimental framework for benchmarking different PHI detection pipelines. We employ three synthetic pixel-level PHI datasets: RadPHI-train for fine-tuning the text localization module, RadPHI-test for evaluating different combinations of YOLOv11, EasyOCR, and GPT-4o, and MIDI for further validation of the optimal setup derived from the RadPHI-test assessments
2 Datasets
For the study at hand, we generate and utilize three distinct datasets: RadPHI-train as our training dataset, and two evaluation datasets, RadPHI-test and MIDI. RadPHI-train and RadPHI-test are created by overlaying synthetically generated imprints on curated medical images across various modalities. MIDI is created by overlaying information stored in DICOM tags, which may contain PHI, utilizing a standard DICOM viewer. The RadPHI-train is used to develop our text detection model, while the RadPHI-test is employed to evaluate different configurations of the PHI detection pipeline. Additionally, the MIDI dataset is reserved as a hold-out test set to assess the optimal configuration identified through the RadPHI-test evaluation. In the upcoming sections, we will elaborate on the generation process for each of these datasets.
2.1 PHI definition
For data simulation and evaluation, we define a list of PHI and non-PHI items based on the HIPAA guidelines [1] as detailed in Table 1. The eight PHI categories include date, general identifier, patient name, address, age, gender, phone number, and email. While the HIPAA guidelines recommend that only ages over 89 are considered PHI, we classify any age as PHI. This is to avoid re-identification, as age groups can exhibit skewed distribution depending on the study, making participants identifiable through age data. Additionally, we group all identification numbers into a group of identifiers, including patient ID, insurance number, social security number, and any number or series that can be used to identify an individual. The eight non-PHI categories include height, weight, examination type, hospital, marker, scanner, diagnosis, and imaging personnel.
Category Description Classification Example
Date All elements of dates related to an individual, including birth date, admission date, discharge date, etc. PHI DOB 01-01-2023
Identifier All identifiers related to an individual such as patient ID, insurance number, medical record number, etc. PHI Patient ID: 0000.0001
Patient name Full name or initials of patient names PHI Pat. Name: John Doe
Address Partial or full address of the patient that contains any of the following: street address, city, state, postal code, and country PHI 123 Main St, Springfield, IL 62701, USA
Age The patient’s age at the time of the scan. PHI Age: 60
Gender The gender of an individual, e.g., male, female, or diverse, or the abbreviations such as m/f/d PHI [M]
Phone number Personal telephone number PHI Contact 794-512-9544
Email Personal email address PHI Email: jane.smith@email.com
Height Height measurement of the patient Non-PHI Height: 165 cm
Weight Weight measurement of the patient Non-PHI Weight 103 kg
Examination type The type of scan performed Non-PHI Exam: CT Cholangiography
Hospital General information about the hospital or imaging facility Non-PHI Mayo Clinic Eau Claire
Marker Text indicating anatomical markers or points of interest Non-PHI R POST L
Scanner General information about the scanner type and scanner settings Non-PHI Philips Ingenia 3.0T
Diagnosis General diagnosis, comments or indications made by doctors Non-PHI Diagnosis: Fibrosis
Imaging personnel Information about the medical staff involved in the scan, such as names of the radiologist, technician, operator, or referring physician who orders, acquires, performs, or interprets the scan Non-PHI Indicated by John Moore
Table 1: Imprint categories and examples. There are 16 categories, eight of which are classified as PHI.
2.2 Preprocessing of public datasets
For the generation of RadPHI-train and RadPHI-test datasets, we curate radiological images from well-known publicly available datasets. Each of them was preprocessed before we added simulated imprints. In the following, we briefly describe each dataset and the applied preprocessing.
Total Segmentator v2
[15] This dataset was originally published and used in a segmentation model for major anatomic structures in CT images. We use the second version, which includes 1,228 CT examinations across 21 scanners. Each volume is resampled to isotropic spacing of 0.45 mm across all three planes. Afterwards, four 2D axial slices are uniformly sampled per volume, which makes a total of 4,896 2D images. The images are min-max normalized to 8-bit format with the minimum value being random between the 0-th and 10-th percentile of the image intensities.
BS-80K
[16] This dataset includes 82,544 radionuclide bone scan images of 13 body regions designed for classification and object detection tasks. Experts de-identified and annotated the images for classification and object detection tasks. For our experiments, we select subsets of whole-body images in anterior and posterior views. Due to the small size of whole-body images, we create image collages of three to five randomly repeated images. This approach mimics a bone scan viewer, where multiple views of a subject are displayed side by side for comparison. To enhance the diversity of bone scan variants, backgrounds are randomly flipped. The processed images are then saved in 8-bit format.
ChestX-ray8
[17] This dataset contains 108,948 frontal X-ray images with nine disease labels parsed from radiological reports. Many images contain physical markers indicating left and right direction. To avoid confusion with added text imprints, we use an object detection model to identify these markers and apply center cropping to remove them. After cropping, the images are saved in 8-bit format.
BRATS
[18] This dataset is part of the Medical Segmentation Decathlon and includes 750 brain MRI images from four sequences: native T1, post-contrast T1, T2-weighted, and T2-FLAIR. All images are resampled to isotropic spacing of 1 mm. For each subject, a random sequence is selected, and four 2D axial slices are uniformly sampled. The images are then min-max normalized to 8-bit format, following a similar approach to the Total Segmentator v2 dataset.
2.3 RadPHI-train
The main use of the RadPHI-train dataset is to enhance the robustness of the text localization module (see Section 3.1) against diverse radiological image backgrounds, varying numbers of imprints, and different imprint representations. This dataset is generated based on images from TotalSegmentator, BS-80K, and ChestX-ray8, applying the preprocessing techniques outlined in Section 2.2. Additionally, we simulate imprints on synthetic backgrounds, incorporating plain grayscale colors along with or without multiple boxes in various shades. The text localization module is trained to manage diverse representations of the imprint rather than focusing on its text content. To achieve this, we continuously simulate the imprints using different fonts, sizes, colors, and locations. These imprints may overlap with each other or with anatomical structures. Although such overlaps are not typically seen in real-world data, we intentionally include them as challenging cases for the model to adapt to. A sample from this dataset is shown in Figure C1 (left) in the Supplementary Material. In total, we generate 6,000 images, which are then divided into training and validation sets using an 80%-20% split.
2.4 RadPHI-test
Figure 3: Imprint simulation pipeline and examples. (Left) Given the input PHI ratio for each dataset, each image can have no imprints, non-PHI imprints, or at least one imprint that contains PHI. The maximum of imprints per image is eight. (Right) Examples of images with imprints across four modalities: (a) Whole body CT (b) Whole body bone scan (c) Chest X-ray (d) Brain MRI.
[…]
Figure 4: Distribution of PHI categories in the RadPHI-test and MIDI datasets. In the RadPHI-test dataset, each image can have a maximum of one imprint per category. In contrast, the MIDI dataset allows for multiple imprints of the same category in each image. We compute the distribution of PHI categories in the MIDI dataset at the image level, meaning that regardless of the number of imprints for a particular category, it is counted as one.
Figure 5: Distribution of PHI and non-PHI imprints in RadPHI-test and MIDI datasets. In RadPHI-test, the number of imprints is limited to eight. In contrast, MIDI has significantly more imprints, correlating to the various DICOM tags randomly displayed on the viewer.
2.5 MIDI
The MIDI dataset is designed to be a challenging hold-out dataset with realistic imprint visualization and placement via a modern DICOM viewer rendering. The image data is curated from the validation set of the 2024 Medical Image De-Identification Benchmark (MIDI-B) challenge [19]. This dataset originally consists of 228 DICOM images across multiple modalities, each containing synthetic PHI content embedded at both the DICOM header and pixel level. We utilize a DICOM viewer, specifically MD.ai [20], to overlay the DICOM tags onto the images. We randomly sample DICOM tags to ensure that the generated imprints represent all possible PHI categories, similar to the RadPHI-test dataset. After applying the overlays, we export the images from the viewer. The resulting images may include not only the DICOM tag overlays but also burn-ins by the challenge organizers. The final version of the dataset comprises 200 images categorized into seven PHI types. The category Email is omitted because emails are neither burned in nor inserted into the header of the original dataset. In contrast to the RadPHI-test dataset, each image can have multiple imprints of the same PHI category. For instance, as illustrated in Figure 6, dates may appear in various contexts, such as the study date, image series, or even the patient comments. We annotate the images at the level of whether any PHI category is present; thus, even if there are numerous imprints of the same category, they are counted as a single instance in our annotation. Figure 4 displays the distribution of PHI categories at the image level. Dates and patient names are among the most frequently occurring PHI elements in the dataset. While we intentionally limit the number of imprints to eight in RadPHI-test, the number of imprints in MIDI can be up to 41, with the majority between 10 and 30 per image (Figure 5).
Figure 6: An example of an image from the MIDI dataset featuring PHI overlays. The same PHI element, such as dates, may appear in various contexts, including patient comments and image series. Besides overlays added by the DICOM viewer (left and right), there are burn-ins by the organizer of the MIDI-B challenge (middle).
3 Methods
3.1 PHI Detection Pipeline
The general design of the PHI detection pipeline comprises three major modules: text localization, text extraction, and text analysis (see Figure 1). The text localization module can be an object detection model that identifies pixels containing text in the input image and returns bounding box coordinates. Once the text coordinates are determined, the text extraction module converts the text within these coordinates into machine-readable text, using OCR algorithms. The machine-encoded text is then analyzed semantically by a language model, which classifies whether it contains any PHI. In the last step, the pipeline flags the image as containing PHI, prompting further review and potential redaction.
Text localization
YOLOv11 [21] is the latest version in the series of YOLO models [22], offering superior performance in real-time object detection tasks. The YOLOv11 architecture is a single-shot object detector consisting of three primary components: a backbone, a neck, and prediction heads. The backbone extracts image features, which are then processed and aggregated across multiple scales in the neck. The final predictions are produced by classification and detection heads. The main improvement in YOLOv11 over the previous version is an optimized backbone and neck, which enhance the extracted features and enable more precise object detection. For our study, we use the small variant of YOLOv11 pre-trained on the COCO dataset and fine-tune it on the RadPHI-train dataset. The model is trained to predict bounding boxes around a single line of text in the image, as illustrated in Figure 1. In addition, we combine real medical images with synthetic images of diverse background colors and shapes to help the model remain robust across all modalities. We use the ultralytics package [21] to train the model with a fixed image size of 640x640 for 100 epochs. The other training arguments are kept as default in the package. Apart from regular geometric augmentation techniques, the images undergo mosaic augmentation, where a new image is created based on patches of original images, as shown in Figure C1 (bottom right). The mosaic augmentation helps the model to generalize better to diverse backgrounds and object locations.
Text extraction
EasyOCR [23] is an open-source OCR package supporting over 80 languages. It comprises two primary modules: recognition and detection. The detection module, analogous to our localization module, identifies text regions in an image using the Character Region Awareness For Text detection (CRAFT) model [24]. The recognition module [25] extracts text by leveraging the Scene Text Recognition module, a multi-stage model that transforms input images, extracts features, aggregates contextual information, and finally outputs the characters. In this study, we use EasyOCR as an out-of-the-box tool without re-training or fine-tuning its components. We keep the values of all parameters as default except for the low-bound text threshold. We reduced this threshold to 0.2 to include more space around detected texts by the CRAFT model, which we found helpful in avoiding missing texts as single letters like markers (e.g., L or R) or abbreviations of gender (e.g., M or F or D). The input to EasyOCR can be either a full image or a cropped image of the text region detected by the localization model. In our experiments, we compare both variants, see Section 4.
Text analysis
GPT-4o [14] is OpenAI’s flagship language model in the Generative Pre-trained Transformer (GPT) family. These models are pre-trained on extensive unlabeled text datasets to generate human-like responses. GPT models have demonstrated their versatility in information extraction and NER tasks, particularly in biomedical text analysis [11]. Leveraging the superior performance of GPT in information extraction, we use GPT-4o to analyze the context of extracted texts from images in order to determine whether they contain PHI and which type of PHI is present. The extracted texts from the text extraction module are provided to GPT-4o as a list. The system prompt consists of the definition of PHI and non-PHI texts and instructions to evaluate the aggregated context of the input text rather than analyzing each line in isolation. Additionally, we provide examples in the prompt to make the model aware of ambiguous cases, following the few-shot prompting technique [26]. For instance, Age: or 65 contains a placeholder for age or hints at a number which could be age, but neither is considered PHI. To ensure consistent outputs, we constrain GPT-4o to produce a dictionary of mandatory fields through OpenAI’s pydantic tool, as outlined in Table 2. If the PHI is detected, the type field specifies one of the eight PHI classes or other if the PHI category is not included in the provided list. If no PHI is detected, the text analyzer assigns non-phi to the type field. Additional fields include raw_text, reason, and language, which store the raw extracted text, the justification for classifying the text as PHI or non-PHI, and the detected language of the original text, respectively. The temperature parameter is set to 0 to yield more deterministic responses.
Tag Data type Description
type str One of [date, identifier, name, address, age, gender, phone_nr, email, other, non-phi]
raw_text str Original text extracted from the image
reason str Reason why the extracted text is considered PHI or not
language str Detected language
Table 2: Structured output of the text analyzer GPT-4o.
4 Experiment
In this study, we assess various setups of a PHI detection pipeline by interchanging the roles of the presented models. Specifically, we study the four setups illustrated in Figure 2. The first setup employs three distinct models, each dedicated to one of the modules described in Section 3.1. This serves as a baseline setup, where each model performs the task it is specifically designed for. In the second setup, we assess the robustness of EasyOCR’s built-in object detection capabilities by using it for both text localization and extraction tasks. EasyOCR handles the detection of text regions and subsequently extracts the corresponding text. The third setup evaluates the language-vision capability of GPT-4o by assigning the text extraction step to it. In this configuration, GPT-4o receives cropped text regions, localized by the YOLOv11 model, performs OCR, analyzes, and classifies PHI content. The final setup explores GPT-4o as an end-to-end solution. GPT-4o takes full responsibility for all three pipeline steps: localizing text in the image, extracting the text, and analyzing it for PHI content. This setup, in particular, highlights a new use case where a single language model integrates vision and language tasks, demonstrating the potential for end-to-end applications.
For each of the introduced setups, we run the pipeline five times on the RadPHI-test dataset to assess the non-deterministic behavior of GPT-4o in generating responses. The final output of the PHI pipeline is evaluated in two stages. In the first stage, the pipeline is treated as a binary classifier for detecting PHI. The objective is to predict whether or not the image contains any PHI. In the second stage, only images identified as containing PHI are considered. The goal here is to evaluate whether the PHI categories are correctly classified. This stage functions as a multi-label classification task across eight PHI classes, see Table 1. In both stages, we calculate the mean values of accuracy, precision, recall, and F1 score by aggregating results from five distinct runs for each setup. The setup that performs best on the RadPHI-test dataset is then further evaluated on the MIDI dataset using the same evaluation metrics.
5 Results
5.1 Rad-PHI dataset
(a) Precision
(b) Recall
(c) F1 Score
(d) Accuracy
Figure 7: Comparison of precision, recall, F1 score, and accuracy across four different setups, averaged from five separate runs of each setup obtained on the RadPHI-test dataset.
Precision Recall F1 Acc
Min Max Min Max Min Max Min Max
Global PHI 0.9875 0.9995 0.9482 0.9972 0.9701 0.9983 0.9506 0.9972
Date 0.9410 1.0000 0.8703 1.0000 0.9307 1.0000 0.9654 1.0000
Patient Name 0.6913 0.9908 0.9025 0.9969 0.7827 0.9938 0.8786 0.9968
Address 0.8725 1.0000 0.9571 0.9992 0.9128 0.9984 0.9563 0.9992
Gender 0.9160 0.9967 0.6934 0.9483 0.8178 0.9714 0.9245 0.9864
Email 0.9813 1.0000 0.8356 0.9964 0.9101 0.9985 0.9542 0.9990
Phone Nr 0.9648 0.9985 0.8339 0.9977 0.9056 0.9981 0.9548 0.9990
Age 0.9246 1.0000 0.6547 0.9913 0.7891 0.9957 0.9103 0.9978
Identifier 0.8026 1.0000 0.8886 0.9797 0.8685 0.9897 0.9359 0.9952
Table 3: Minimum and maximum scores of mean precision, mean recall, mean F1, and mean accuracy for each PHI category computed across all setups obtained on RadPHI-test. The color-coding indicates the setup attaining the corresponding metric value, i.e. Setup 1, Setup 2, Setup 3, Setup 4.
Setup Models Pipeline error rate Total time Prompt tokens Response tokens
1 YOLOv11 + EasyOCR + GPT-4o $0.080.08$ % 3,64136413,6413 , 641 s 1,686,27716862771,686,2771 , 686 , 277 227,702227702227,702227 , 702
2 EasyOCR + GPT-4o $0.100.10$ % 4,78547854,7854 , 785 s 1,704,06217040621,704,0621 , 704 , 062 259,819259819259,819259 , 819
3 YOLOv11 + GPT-4o $0.280.28$ % 12,2951229512,29512 , 295 s 5,639,64956396495,639,6495 , 639 , 649 425,935425935425,935425 , 935
4 GPT-4o $8.568.56$ % 3,43434343,$43$ , 434 s 3,249,45532494553,249,4553 , 249 , 455 250,577250577250,577250 , 577
Table 4: Comparison of mean error rate, mean total time of analysis, the mean number of prompt and response tokens for 1,000 test cases across five runs for each setup evaluated on RadPHI-test. The pipeline error rate is the percentage of images that fail to process due to server issues or the presence of banned words in the requests sent to GPT-4o. The total time is estimated as the sum of the processing and analysis time required for the entire dataset.
Figure 8: When text appears in multi-line blocks, out-of-the-box EasyOCR tries to fit bounding boxes that cover the entire text space. However, the bounding boxes might break the context consistency of the text. For example, Visit Date and 03-01-1989 are separated into two different crops, making it harder for the text analyzer to classify PHI content, particularly when a more fine-grain classification of dates is required, e.g., visit date or date of birth.
The results of the four setups are summarized in Figure 7, showing the precision, recall, F1 score, and accuracy in classifying images for global PHI (i.e., whether the image has PHI or not) and the corresponding eight PHI categories. Table 3 compares the pipeline error rate, total time taken to process 1000 images, and the number of prompt and response tokens used by GPT-4o across four setups. The pipeline error rate indicates the percentage of images that couldn’t complete the process due to GPT-4o’s inability to handle the request or the presence of its banned words in the query. Overall, all four setups demonstrate strong performance in deciding if an image contains PHI imprints, with all evaluation metrics related to the Global PHI label exceeding $0.940.94$. Setup 2 (EasyOCR + GPT-4o) and setup 4 (GPT-4o) exhibit slightly lower performance compared to the other setups. However, their performance remains robust, with the overall accuracy at $0.96040.9604$ and $0.95060.9506$, respectively. Setup 3 (YOLOv11 + GPT-4o) achieves the highest performance, with all four metrics surpassing $0.990.99$. This indicates that the combination of YOLOv11 for text localization and GPT-4o for text extraction and analysis excels at identifying images containing PHI. The overall high scores in all evaluation metrics, even in lower-performing setups, suggest that detecting the presence of PHI containing imprints is for all pipeline setups a solvable task.
Assessing the capability of different pipeline setups to perform fine-grained classification of PHI categories reveals strong performance deviations between the various setups. First, we observe that the PHI classes, such as date and address, are among the top-performing categories across all setups. This result is expected because both categories exhibit distinct semantic structures compared to other classes. On the other hand, gender information is the most challenging PHI class to detect, in particular when this information appears in abbreviated forms. Single-letter gender abbreviations often lack accompanying text like gender or sex, making detection more difficult. If the text extractor fails to recognize the exact character, the text analyzer is left without contextual clues to identify the PHI content and class. This issue is further complicated by potential confusion with single-character markers like L, R, A or P. Generally speaking, the bottleneck for fine-grained PHI category classification generally occurs during the text extraction step. When texts are incompletely extracted or when signaling texts are missed, it becomes significantly harder for the text analyzer to classify the content accurately. For example, Figure 8 illustrates how EasyOCR can break text continuity into separate bounding boxes, resulting in missing context and complicating the advanced analysis of PHI terms.
Among the four setups, Setup 2 has one of the lowest performance levels. This outcome is expected, as the detection component of EasyOCR is not trained on medical image data, making it less robust compared to a dedicated YOLOv11 model. Even when EasyOCR is used solely for text extraction (Setup 1), the final pipeline performance lags slightly behind Setup 3 where GPT-4o is used for integrated text extraction and analysis. Upon manual examination of specific cases, EasyOCR demonstrates more errors in text extraction compared to using GPT-4o for this task. However, in a majority of cases, GPT-4o, when used as a text analyzer, is powerful enough to infer the intended meaning from improperly extracted text by EasyOCR. This capability results in a comparable performance between Setup 1 and 3. While Setup 3 offers a slightly better overall performance, it has a significant drawback: the total analysis time is more three times that of Setup 1. This increase is primarily due to the latency associated with querying GPT-4o multiple times for extracting text from individual image crops. On average, the number of prompt and response tokens in Setup 3 is three times and two times higher than that in Setup 1, as shown in Table 4. Using GPT-4o for the entire pipeline is technically feasible, but it does carry a higher rate of pipeline errors and hallucinations. These errors frequently arise from unstructured outputs or the generation of special tokens that can cause the model to fail to respond. While our experimental result underscores the impressive vision capabilities of GPT-4o, it is still inadequate to entirely replace specialized models for text localization and extraction tasks in a single pass. For further evaluation on the MIDI dataset, we choose Setup 1 over Setup 3 as the gain in performance of the latter is very limited whereas inference costs, latency and token costs are much more favorable for Setup 1.
5.2 MIDI
In comparison to our simulated RadPHI-test dataset, the MIDI dataset is more challenging due to its increased diversity and a larger number of imprints. Relying on signaling words might not be sufficient to detect PHI effectively. For example, an image series by itself may not be classified PHI, but if it includes a date element, as shown in Figure 6, it is considered PHI. Overall, the performance of Setup 1 on MIDI is comparable to that obtained on RadPHI-test, see Figure 9 (left). Notably, the global PHI detection has a recall rate close to 1. Key categories demonstrating strong performance include dates, patient names, and addresses. However, similar to findings from the RadPHI-test, the identification of gender remains particularly challenging, especially in the context of abbreviations. The precision of the identifier category is diminished due to misclassifications of numerous image- and study-related identifiers as PHI. These misclassifications arise from our initial prompt to GPT-4o, which states that all forms of identifiers are regarded as PHI. To address the needs of this scenario, we refine the prompt by explicitly stating that study or image-related identifiers should not be classified as PHI. The improvement in precision for the identifier category following this prompt adjustment is demonstrated in Figures 9 (right) and  10. This case serves as a compelling illustration of how LLMs can adapt to detect PHI based on specific criteria, rather than being confined to a fixed definition. Overall, this allows users to configure the PHI detection system more flexibly, for example, to meet study-specific criteria.
(a) Same system prompt as in RadPHI-test
(b) Modified system prompt for identifier
Figure 9: Performance comparison of precision, recall, F1 score, and accuracy on the MIDI dataset using Setup 1: (a) performance with the RadPHI-test prompt and (b) performance with an adapted prompt for identifier classification.
Figure 10: (Left) DICOM tag overlays with the DICOM viewer. (Right) Pipeline output for Setup 1 on the MIDI dataset: (a) using a prompt that considers all identifiers as PHI, and (b) with a modified prompt excluding study/image identifiers. Despite SERIES being non-PHI in the new prompt, the LLM can still detect the date element.
6 Discussion
We investigate four different setups of an end-to-end pipeline for pixel-level PHI detection in medical images. For the localization step, we recommend using a dedicated object detection model instead of relying on the built-in localization component of a generic OCR model (Setup 2) or the localization capability of a multi-modal LLM (Setup 4). Our experimental results demonstrate that separating the localization step and using a localization model trained explicitly on the data reflecting the target domain ensures strong overall performance, even when out-of-the-box OCR tools, like EasyOCR, are used downstream for text extraction (Setup 1). End-to-end GPT-4o (Setup 4) has the drawback of not returning the bounding box coordinates of text regions, which is essential for subsequent PHI redaction. Additionally, imprints are often incompletely identified and hallucinations occur much more frequently compared to other setups, resulting in lower performance. Moreover, there are more pipeline errors related to server issues or unstructured output, which prevents GPT-4o from generating a proper response.
For the text extraction step, we observe that GPT-4o outperforms EasyOCR by having lower word-level and character-level errors. However, when evaluating the end-to-end pipeline performance, the improvement in PHI detection when switching to GPT-4o from EasyOCR (i.e, Setup 3 vs. Setup 1) is minimal. This is because minor OCR errors produced by EasyOCR can easily be circumvented by the robust language model in the text analysis step. While GPT-4o offers superior performance in text extraction, the overall gain in pipeline performance is modest and accompanied by higher latency and increased costs associated with token generation. Therefore, EasyOCR is preferred in our study.
For text analysis, GPT-4o is the default and only choice across four different setups. The model excels at understanding text content and is particularly efficient in identifying PHI and further classifying it into sub-categories. Additionally, it can accurately discern context even when there are slight OCR errors in the extracted text. For instance, Figure 11 illustrates OCR spelling errors related to dates, identifiers, and gender. The model is capable of implicitly correcting these errors and reasoning based on the corrected text. Furthermore, one of the strengths of LLM is its flexibility, which allows for adjustments in instructions through prompts and configurations of output structures. This adaptability makes it easy to redefine PHI and associated tasks across various studies.
Figure 11: (Left) An image from the RadPHI-test dataset with simulated imprints. (Right) Results from the pipeline utilizing EasyOCR for text extraction and GPT-4o for text analysis. Despite the presence of spelling errors in the extracted text, GPT-4o effectively recognizes and interprets the content.
A key concern for any AI-based system is its ability to generalize to unseen data. In light of this, we explore the use of pre-trained or foundational models with minimal fine-tuning, applicable solely to the localization model. Evaluation on the hold-out MIDI dataset has demonstrated that the YOLO model, despite being fine-tuned on a different dataset (RadPHI-train), can perform effectively on new datasets containing modalities not encountered during training, including PET, mammography, and ultrasound. Furthermore, the LLM demonstrates adaptability to diverse contexts, as evidenced by our results on the MIDI dataset with modified prompting. This underscores a valuable use case in which users can interactively adjust the prompt to redact or retain specific information in the images, all without the need to develop a new classification model. Overall, these findings emphasize the potential of leveraging existing models to enhance performance across various medical imaging domains.
In terms of limitations, the expansion of GPT-4o’s capabilities beyond basic text analysis to the interpretation of small image crops and, ultimately, entire images reveals interesting findings. While the model demonstrates exceptional performance when analyzing image crops (Setup 3), it encounters a significant increase in technical errors and hallucinations when processing complete images (Table 3). This inefficiency results in incomplete and inaccurate detection of PHI. It also underscores the current limitations of the multimodal language model, particularly in its capacity to handle images with diverse content and to perform complex reasoning tasks that require the integration of information from both images and text. Moreover, the model’s non-deterministic behavior can lead to varying responses for identical inputs. However, by employing structured output formats with Pydantic [27] and setting the temperature parameter to zero, we have observed a marked improvement in response consistency, as indicated by the low standard deviation (Table A1-A4 in Supplementary Material). There remains significant potential to further explore the full capabilities of language models for the PHI detection task. In our future work, we plan to investigate and benchmark various prompting strategies and assess the model’s robustness in languages other than English. Additionally, given the high level of privacy required for medical data, it is essential to either fine-tune a LLM or host it in-house for application development. Smaller LLMs may be particularly interesting to examine due to their ease of deployment and the ability to customize them for specific tasks. Furthermore, we plan to benchmark our pipeline against commercially available products [8, 7, 9, 10] currently available in the market.
7 Conclusion
In this paper, we dissect a pixel-level PHI detection pipeline into fundamental components which are implemented using well-established and generic foundation models. In detail, we assess various pipeline setups utilizing YOLOv11 as an object detection model, EasyOCR as an OCR model, and GPT-4o as a text analysis model. Our findings indicate that employing LLM for the analysis phase yields remarkable contextual understanding and offers the flexibility to configuration adjustment. For the text localization and extraction phases, the optimal approach combines a fine-tuned object detection model with an out-of-the-box OCR tool. This combination outperforms pipelines relying on the OCR model’s built-in text detection. Notably, we demonstrate that the LLM can extract text from image crops with performance levels that are on par with, or even exceed, those of established OCR tools like EasyOCR. However, while utilizing GPT-4o as an end-to-end PHI detection solution is a feasible approach, the setups relying on dedicated vision models designed for comprehensive image analysis are still superior in this case. This gap points to a promising area for future research aimed at enhancing vision-integrated LLM performance.
Appendix A Results on RadPHI-test
Class Precision Recall F1 Accuracy
Global PHI 0.9988 ± 0.0000 0.9904 ± 0.0016 0.9946 ± 0.0008 0.9908 ± 0.0013
Date 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000
Patient Name 0.9493 ± 0.0030 0.9961 ± 0.0001 0.9722 ± 0.0016 0.9852 ± 0.0008
Address 0.9976 ± 0.0022 0.9992 ± 0.0018 0.9984 ± 0.0017 0.9992 ± 0.0008
Gender 0.9967 ± 0.0030 0.7414 ± 0.0062 0.8512 ± 0.0059 0.9363 ± 0.0015
Email 1.0000 ± 0.0000 0.9509 ± 0.0033 0.9749 ± 0.0017 0.9864 ± 0.0009
Phone Nr 0.9923 ± 0.0000 0.9923 ± 0.0000 0.9923 ± 0.0000 0.9960 ± 0.0000
Age 1.0000 ± 0.0000 0.9266 ± 0.0051 0.9619 ± 0.0028 0.9812 ± 0.0013
Identifier 0.9948 ± 0.0019 0.9747 ± 0.0030 0.9847 ± 0.0018 0.9928 ± 0.0008
Table A1: PHI detection performance of setup 1 (YOLOv11 + EasyOCR + GPT-4o) on RadPHI-test.
Class Precision Recall F1 Accuracy
Global PHI 0.9875 ± 0.0011 0.9656 ± 0.0013 0.9764 ± 0.0009 0.9604 ± 0.0016
Date 1.0000 ± 0.0000 0.8703 ± 0.0063 0.9307 ± 0.0037 0.9654 ± 0.0017
Patient Name 0.8480 ± 0.0039 0.9308 ± 0.0027 0.8874 ± 0.0023 0.9385 ± 0.0014
Address 0.9812 ± 0.0022 0.9694 ± 0.0036 0.9752 ± 0.0023 0.9878 ± 0.0011
Gender 0.9965 ± 0.0032 0.6934 ± 0.0034 0.8178 ± 0.0029 0.9245 ± 0.0012
Email 0.9991 ± 0.0019 0.8356 ± 0.0042 0.9101 ± 0.0031 0.9542 ± 0.0015
Phone Nr 0.9909 ± 0.0032 0.8339 ± 0.0032 0.9056 ± 0.0027 0.9548 ± 0.0013
Age 0.9929 ± 0.0026 0.6547 ± 0.0045 0.7891 ± 0.0032 0.9103 ± 0.0010
Identifier 0.8492 ± 0.0061 0.8886 ± 0.0056 0.8685 ± 0.0038 0.9361 ± 0.0019
Table A2: PHI detection performance of setup 2 (EasyOCR + GPT-4o) on RadPHI-test.
Class Precision Recall F1 Accuracy
Global PHI 0.9995 ± 0.0007 0.9972 ± 0.0006 0.9983 ± 0.0005 0.9972 ± 0.0008
Date 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000
Patient Name 0.9908 ± 0.0021 0.9969 ± 0.0017 0.9938 ± 0.0008 0.9968 ± 0.0004
Address 1.0000 ± 0.0000 0.9959 ± 0.0001 0.9980 ± 0.0000 0.9990 ± 0.0000
Gender 0.9957 ± 0.0000 0.9483 ± 0.0063 0.9714 ± 0.0033 0.9864 ± 0.0016
Email 1.0000 ± 0.0000 0.9964 ± 0.0000 0.9985 ± 0.0007 0.9990 ± 0.0000
Phone Nr 0.9985 ± 0.0021 0.9977 ± 0.0021 0.9981 ± 0.0014 0.9990 ± 0.0007
Age 1.0000 ± 0.0000 0.9913 ± 0.0018 0.9957 ± 0.0009 0.9978 ± 0.0004
Identifier 1.0000 ± 0.0000 0.9797 ± 0.0082 0.9897 ± 0.0042 0.9952 ± 0.0019
Table A3: PHI detection performance of setup 3 (YOLOv11 + GPT-4o) on RadPHI-test.
Class Precision Recall F1 Accuracy
Global PHI 0.9929 ± 0.0015 0.9482 ± 0.0030 0.9701 ± 0.0013 0.9506 ± 0.0021
Date 0.9410 ± 0.0072 0.9513 ± 0.0050 0.9461 ± 0.0030 0.9722 ± 0.0017
Patient Name 0.6913 ± 0.0147 0.9025 ± 0.0146 0.7827 ± 0.0082 0.8786 ± 0.0063
Address 0.8725 ± 0.0048 0.9571 ± 0.0083 0.9128 ± 0.0039 0.9563 ± 0.0021
Gender 0.9160 ± 0.0090 0.7560 ± 0.0095 0.8284 ± 0.0081 0.9278 ± 0.0024
Email 0.9813 ± 0.0019 0.9034 ± 0.0075 0.9407 ± 0.0038 0.9696 ± 0.0017
Phone Nr 0.9648 ± 0.0063 0.8697 ± 0.0117 0.9148 ± 0.0068 0.9597 ± 0.0031
Age 0.9246 ± 0.0089 0.9431 ± 0.0067 0.9337 ± 0.0038 0.9666 ± 0.0022
Identifier 0.8026 ± 0.0109 0.9550 ± 0.0073 0.8722 ± 0.0084 0.9359 ± 0.0050
Table A4: PHI detection performance of setup 4 (GPT-4o) on RadPHI-test.
Appendix B Results on MIDI
Class Precision Recall F1 Accuracy
Global PHI 0.9900 ± 0.0034 0.9962 ± 0.0035 0.9931 ± 0.0014 0.9890 ± 0.0022
Date 1.0000 ± 0.0000 0.9948 ± 0.0077 0.9974 ± 0.0039 0.9970 ± 0.0045
Patient Name 0.9735 ± 0.0076 0.9885 ± 0.0105 0.9809 ± 0.0059 0.9800 ± 0.0061
Address 0.9631 ± 0.0291 0.9889 ± 0.0152 0.9755 ± 0.0110 0.9910 ± 0.0042
Gender 0.9965 ± 0.0079 0.9492 ± 0.0000 0.9722 ± 0.0038 0.9840 ± 0.0022
Phone Nr 0.9723 ± 0.0253 0.9619 ± 0.0399 0.9663 ± 0.0138 0.9930 ± 0.0027
Age 0.9637 ± 0.0133 0.9606 ± 0.0083 0.9621 ± 0.0075 0.9750 ± 0.0050
Identifier 0.7622 ± 0.0093 0.9899 ± 0.0106 0.8612 ± 0.0088 0.8740 ± 0.0082
Table B1: PHI detection performance of setup 1 (YOLOv11 + EasyOCR + GPT-4o) on MIDI using the same prompt as RadPHI-test
Class Precision Recall F1 Accuracy
Global PHI 0.9987 ± 0.0028 0.9924 ± 0.0029 0.9955 ± 0.0017 0.9930 ± 0.0027
Date 0.9983 ± 0.0038 0.9931 ± 0.0038 0.9957 ± 0.0030 0.9950 ± 0.0035
Patient Name 0.9923 ± 0.0124 0.9808 ± 0.0000 0.9865 ± 0.0062 0.9860 ± 0.0065
Address 0.9472 ± 0.0255 0.9889 ± 0.0152 0.9675 ± 0.0181 0.9880 ± 0.0067
Gender 0.9965 ± 0.0078 0.9627 ± 0.0142 0.9793 ± 0.0099 0.9880 ± 0.0057
Phone Nr 0.9735 ± 0.0391 1.0000 ± 0.0000 0.9862 ± 0.0204 0.9970 ± 0.0045
Age 0.9639 ± 0.0079 0.9697 ± 0.0000 0.9668 ± 0.0040 0.9780 ± 0.0027
Identifier 0.9920 ± 0.0118 0.9342 ± 0.0165 0.9621 ± 0.0087 0.9709 ± 0.0066
Table B2: PHI detection performance of setup 1 (YOLOv11 + EasyOCR + GPT-4o) on MIDI using the modified prompt with focus on eliminating false positives in the identifier category.
Appendix C Training data for YOLOv11
Figure C1: Examples of data used for training the localization model YOLOv11. (Left) Training image with multiple simulated text imprints in random locations (Right) Augmentation of training images. One of the augmentation schemes we use is mosaic augmentation, where the new image is created by blending multiple patches of original images.
Appendix D Examples of pipeline output
Figure D1: Comparison of outputs from four setups: Setups 1, 3, and 4 demonstrate similar response patterns, with Setup 3 accurately identifying the gender. In contrast, Setup 2 exhibits incomplete text extraction, missing the patient’s name and gender.
Figure D2: Comparison of outputs from four setups: Setup 1 and Setup 3 accurately classify PHI categories despite overlapping address image regions. Setup 2 fails to recognize the age class, while Setup 4 illustrates model hallucination, generating responses that are not based on the provided input.
Figure D3: Comparison of responses from the same case across five distinct runs. Utilizing Pydantic and a temperature setting of 0 enhances the consistency of responses, even though some variation in reasoning is observed across the runs.

### v4

@@
 Despite the importance of such systems, there has been limited evaluation of existing AI-based solutions, creating barriers to the development of reliable and robust tools.
-In this study, we present an AI-based pipeline for PHI detection, comprising three key components: text detection, text extraction, and text analysis.
-We benchmark three models—YOLOv11, EasyOCR, and GPT-4o— across different setups corresponding to these components, evaluating the performance based precision, recall, F1 score, and accuracy.
-All setups demonstrate excellent PHI detection, with all metrics exceeding 0.9.
-The combination of YOLOv11 for text localization and GPT-4o for extraction and analysis yields the best results.
-However, this setup incurs higher costs due to GPT-4o’s token generation.
-Conversely, an end-to-end pipeline that relies solely on GPT-4o shows lower performance but highlights the potential of multi-modal models for complex tasks.
-We recommend fine-tuning a dedicated object detection model and utilizing built-in OCR tools to achieve optimal performance and cost-effectiveness.
-Additionally, leveraging language models such as GPT-4o can facilitate thorough and flexible analysis of text content.
+In this study, we present an AI-based pipeline for PHI detection, comprising three key modules: text detection, text extraction, and text analysis.
+We benchmark three models—YOLOv11, EasyOCR, and GPT-4o— across different setups corresponding to these modules, evaluating their performance on two different datasets encompassing multiple imaging modalities and PHI categories.
+Our findings indicate that the optimal setup involves utilizing dedicated vision and language models for each module, which achieves a commendable balance in performance, latency, and cost associated with the usage of Large Language Models (LLMs).
+Additionally, we show that the application of LLMs not only involves identifying PHI content but also enhances OCR tasks and facilitates an end-to-end PHI detection pipeline, showcasing promising outcomes through our analysis.
 Keywords: de-identification, Protected Health Information (PHI) detection, medical imaging, Large Language Model (LLM)
@@
 Before sharing data, e.g., for research purposes, it is crucial to de-identify or pseudonymize it to minimize the risk of re-identification.
-This de-identification step is mandatory for studies to comply with the Health Insurance Portability and Accountability Act (HIPAA) [1] and the General Data Protection Regulation (GDPR) [2], and it is the responsibility of the data provider.
+This de-identification step is mandatory for studies to comply with the Health Insurance Portability and Accountability Act (HIPAA) [1] and the General Data Protection Regulation (GDPR) [2], and falls under the responsibility of the data processor.
+Furthermore, fully de-identified datasets enable broader and unrestricted use, facilitating collaborative research and benchmarking without the need for complex and time-consuming processes to obtain patient authorization and approval.
 De-identification is a complex process that involves many data types and formats, typically including a detection step followed by the redaction of the identified PHI.
 Medical images are a rich source of PHI, which can be embedded as metadata (e.g., in DICOM headers) or burned-in text within pixel data.
-The automatic detection of pixel-level PHI is challenging due to the diversity in imprint content, locations, font sizes, color schemes, annotations, and machine overlays.
+However, manually detecting PHI in medical images is infeasible for large datasets, especially those that are curated for machine learning applications.
+Thus, the development of an automated tool for recognizing PHI is not just advantageous but essential for maintaining compliance and ensuring the privacy and security of sensitive health information.
+By leveraging automation, researchers and healthcare professionals can streamline the de-identification process, safeguarding patient privacy while enabling valuable insights from healthcare data.
+Nevertheless, the automatic detection of pixel-level PHI is challenging due to the diversity in imprint content, locations, font sizes, color schemes, annotations, and machine overlays.
 As the overall volume of medical data increases, so does the risk of incomplete PHI redaction, potentially leading to data breaches, which imply significant negative consequences for affected individuals and the healthcare sector, as well as high costs for responsible entities [3].
@@
 Previous works have employed Optical Character Recognition (OCR) technology to recognize text in medical images, leading to two primary approaches for handling burned-in text.
-The first approach removes all burned-in texts, regardless of content [5].
+The first approach leverages text detection or OCR models such as Tesseract [5] to localize burned-in texts and removes all of them, regardless of content [6, 7, 8].
 While this method effectively eliminates any potential PHI, it also discards valuable information, such as measurements and markers.
-The second approach applies heuristic rules to redact specific PHI imprints [6].
-However, this method is tailored to particular representations of imprints in images and can struggle to generalize to new types of PHI or different representations.
+The second approach applies OCR models to extract burned-in texts and designs heuristic rules to recognize PHI based on matching the extracted text to a defined PHI dictionary and subsequently redact them [9, 10, 11].
+While effective in structured settings, rule-based approaches are often tailored to specific imprint formats or language patterns and may struggle to generalize when faced with heterogeneous texts, ambiguous phrasing, or languages not covered by predefined rules.
 In addition, fine-grained classification of PHI is often preferred to coarse classification, which categorizes information simply as names or dates.
-For instance, not all patient identifiers or dates need redaction.
-Identifiers linked to the hospital system pose a high risk of identification and should indeed be redacted, while those used in clinical trials are valid for study analysis.
-Similarly, although all dates are classified as PHI under HIPAA guidelines, certain dates, such as those documenting patient visits, are critical for clinical trials.
+This is particularly needed in certain research contexts where information such as visit dates or study-specific identifiers may be retained under a data use agreement, as it is essential for study analysis.
 As the volume and diversity of medical data continue to grow, advanced automatic solutions for PHI detection are essential, particularly for managing large datasets collected retrospectively.
-These solutions are expected to handle edge cases, which can involve various signaling words and phrases.
-Moreover, research sites require flexibility to determine which entities to include or exclude as PHI based on specific study setup.
-Therefore, a robust PHI detection tool must incorporate mechanisms that enable easy specification and adjustment of rules to meet the needs of different studies.
+Therefore, a robust PHI detection tool must incorporate mechanisms that enable easy specification and adjustment of rules to meet the needs of different studies or country-specific guidelines.
 Our study focuses on AI-based approaches for automatically detecting PHI as burned-in pixels.
-Currently, there are a few solutions that utilize vision and language models from MD.ai [7], Google [8], John Snow Labs [9], and Glendor [10].
+Current commercial tools that address this need include MD.ai [12], Google [13], John Snow Labs [14], and Glendor [15].
+Additionally, Microsoft provides an open-source tool called Presidio [16] that serves similar purposes.
 These tools commonly integrate two key components: a vision component that detects and converts burned-in texts into machine-encoded strings, and a language component that identifies any PHI content in the extracted texts [17].
@@
   • We propose an end-to-end PHI detection system comprising three modules: text localization to identify text in images, text extraction to convert pixels into machine-readable text, and text analysis to detect PHI content using advanced language models (Figure 1).
-  • By leveraging publicly available medical imaging data and the HIPAA guidelines, we create two evaluation datasets comprising 1000 and 200 images, respectively.
-These datasets encompass various radiology modalities, including X-Ray, CT, PET, MRI, bone scintigraphy, mammography, and ultrasound.
-The images are synthesized with pixel-level PHI to support development and benchmarking within the research community.
   • We combine three models (i.e., YOLOv11, EasyOCR, and GPT-4o) corresponding to three previously mentioned modules in four different setups of a PHI detection pipeline (Figure 2) and compare their performance in an extensive quantitative study.
@@
 For example, EasyOCR can perform not only text extraction but also text localization, and GPT-4o can perform end-to-end PHI detection instead of only analyzing the semantic content of extracted texts.
-  • Based on the quantitative results, we provide insights into the optimal pipeline setup, considering performance, latency, and inference cost in terms of the number of generated tokens.
+  • By leveraging publicly available medical imaging data and the HIPAA guidelines, we create two benchmark datasets comprising 1000 and 200 images, respectively.
+These datasets encompass various radiology modalities, including X-Ray, CT, PET, MRI, bone scintigraphy, mammography, and ultrasound.
+For both datasets we provide detailed annotations to support development and benchmarking within the research community.
+  • Based on an extensive quantitative analysis, we provide insights into the optimal pipeline setup, considering performance, latency, and inference cost in terms of the number of generated tokens.
 Specifically, we explore the full capabilities of GPT-4o for both vision and language tasks, and analyze the trade-offs involved in utilizing it partially or entirely in the PHI detection pipeline.
@@
 RadPHI-train and RadPHI-test are created by overlaying synthetically generated imprints on curated medical images across various modalities.
-MIDI is created by overlaying information stored in DICOM tags, which may contain PHI, utilizing a standard DICOM viewer.
+MIDI is created by overlaying information stored in DICOM tags, which may contain PHI, utilizing an industry standard DICOM viewer.
 The RadPHI-train is used to develop our text detection model, while the RadPHI-test is employed to evaluate different configurations of the PHI detection pipeline.
@@
 For data simulation and evaluation, we define a list of PHI and non-PHI items based on the HIPAA guidelines [1] as detailed in Table 1.
-The eight PHI categories include date, general identifier, patient name, address, age, gender, phone number, and email.
-While the HIPAA guidelines recommend that only ages over 89 are considered PHI, we classify any age as PHI.
-This is to avoid re-identification, as age groups can exhibit skewed distribution depending on the study, making participants identifiable through age data.
-Additionally, we group all identification numbers into a group of identifiers, including patient ID, insurance number, social security number, and any number or series that can be used to identify an individual.
-The eight non-PHI categories include height, weight, examination type, hospital, marker, scanner, diagnosis, and imaging personnel.
+The six PHI categories include date, general identifier, patient name, address, phone number, and email.
+We group all identification numbers into a group of identifiers, including patient ID, insurance number, social security number, and any number or series that can be used to identify an individual.
+The ten non-PHI categories include age < 90, gender, height, weight, examination type, hospital, marker, scanner, diagnosis, and imaging personnel.
 Category Description Classification Example
-Date All elements of dates related to an individual, including birth date, admission date, discharge date, etc. PHI DOB 01-01-2023
-Identifier All identifiers related to an individual such as patient ID, insurance number, medical record number, etc. PHI Patient ID: 0000.0001
+Date All elements of dates related to an individual, including birth date, admission date, discharge date, and all ages over 89 years, as well as all elements of dates indicative of such age PHI DOB 01-01-2023
+Identifier All identifiers related to an individual, such as patient identifier, insurance number, and medical record number PHI Patient ID: 0000.0001
 Patient name Full name or initials of patient names PHI Pat. Name: John Doe
 Address Partial or full address of the patient that contains any of the following: street address, city, state, postal code, and country PHI 123 Main St, Springfield, IL 62701, USA
-Age The patient’s age at the time of the scan. PHI Age: 60
-Gender The gender of an individual, e.g., male, female, or diverse, or the abbreviations such as m/f/d PHI [M]
 Phone number Personal telephone number PHI Contact 794-512-9544
 Email Personal email address PHI Email: jane.smith@email.com
+Age < 90 Age that is below 90, as well as all elements of dates indicative of such age. Non-PHI Age: 60
+Gender The gender of an individual, e.g., male, female, or diverse, or the abbreviations such as m/f/d Non-PHI [M]
 Height Height measurement of the patient Non-PHI Height: 165 cm
@@
 Imaging personnel Information about the medical staff involved in the scan, such as names of the radiologist, technician, operator, or referring physician who orders, acquires, performs, or interprets the scan Non-PHI Indicated by John Moore
-Table 1: Imprint categories and examples. There are 16 categories, eight of which are classified as PHI.
+Table 1: Imprint categories and examples. There are 16 categories, six of which are classified as PHI.
 2.2 Preprocessing of public datasets
@@
 2.3 RadPHI-train
-The main use of the RadPHI-train dataset is to enhance the robustness of the text localization module (see Section 3.1) against diverse radiological image backgrounds, varying numbers of imprints, and different imprint representations.
+The main use of the RadPHI-train dataset is to enhance the robustness of the text localization module (see Section 3) against diverse radiological image backgrounds, varying numbers of imprints, and different imprint representations.
 This dataset is generated based on images from TotalSegmentator, BS-80K, and ChestX-ray8, applying the preprocessing techniques outlined in Section 2.2.
@@
 In RadPHI-test, the number of imprints is limited to eight.
-In contrast, MIDI has significantly more imprints, correlating to the various DICOM tags randomly displayed on the viewer.
+In contrast, MIDI has significantly more imprints, correlating to the various DICOM tags and modality-specific imprints randomly displayed on the viewer.
 2.5 MIDI
@@
 The resulting images may include not only the DICOM tag overlays but also burn-ins by the challenge organizers.
-The final version of the dataset comprises 200 images categorized into seven PHI types.
+The final version of the dataset comprises 200 images categorized into five PHI types.
 The category Email is omitted because emails are neither burned in nor inserted into the header of the original dataset.
@@
 For instance, as illustrated in Figure 6, dates may appear in various contexts, such as the study date, image series, or even the patient comments.
-We annotate the images at the level of whether any PHI category is present; thus, even if there are numerous imprints of the same category, they are counted as a single instance in our annotation.
+We performed instance-level annotation of the images by generating coordinates for PHI instances along with their corresponding categories.
+This annotation process was carried out and validated by two independent annotators to ensure accuracy and reliability.
 Figure 4 displays the distribution of PHI categories at the image level.
@@
 3 Methods
-3.1 PHI Detection Pipeline
 The general design of the PHI detection pipeline comprises three major modules: text localization, text extraction, and text analysis (see Figure 1).
@@
 In the last step, the pipeline flags the image as containing PHI, prompting further review and potential redaction.
-Text localization
+In our study, we explore three distinct models: YOLOv11, EasyOCR, and GPT-4o.
+These models are configured in various combinations to fulfill the roles of the three modules mentioned above.
+We will describe each model in detail and outline how they are integrated to perform the functions of text localization, text extraction, and text analysis.
+YOLOv11
 YOLOv11 [27] is the latest version in the series of YOLO models [28], offering superior performance in real-time object detection tasks.
+In our study, we employ YOLOv11 exclusively for the text localization module.
 The YOLOv11 architecture is a single-shot object detector consisting of three primary components: a backbone, a neck, and prediction heads.
@@
 In addition, we combine real medical images with synthetic images of diverse background colors and shapes to help the model remain robust across all modalities.
-We use the ultralytics package [21] to train the model with a fixed image size of 640x640 for 100 epochs.
-The other training arguments are kept as default in the package.
+We use the ultralytics package [27] to fine-tune the model with a fixed image size of 640x640 for 100 epochs.
+The other training hyperparameters are kept as default in the package.
 Apart from regular geometric augmentation techniques, the images undergo mosaic augmentation, where a new image is created based on patches of original images, as shown in Figure C1 (bottom right).
 The mosaic augmentation helps the model to generalize better to diverse backgrounds and object locations.
-Text extraction
+EasyOCR
 EasyOCR [29] is an open-source OCR package supporting over 80 languages.
@@
 The recognition module [31] extracts text by leveraging the Scene Text Recognition module, a multi-stage model that transforms input images, extracts features, aggregates contextual information, and finally outputs the characters.
-In this study, we use EasyOCR as an out-of-the-box tool without re-training or fine-tuning its components.
+We investigate EasyOCR under two distinct setups: (1) as a dedicated text extractor, and (2) as both a text localizer and a text extractor.
+For both setups, we use EasyOCR as an out-of-the-box tool without re-training or fine-tuning its components.
 We keep the values of all parameters as default except for the low-bound text threshold.
 We reduced this threshold to 0.2 to include more space around detected texts by the CRAFT model, which we found helpful in avoiding missing texts as single letters like markers (e.g., L or R) or abbreviations of gender (e.g., M or F or D).
-The input to EasyOCR can be either a full image or a cropped image of the text region detected by the localization model.
-In our experiments, we compare both variants, see Section 4.
-Text analysis
+The input to EasyOCR can be either a full image or a cropped image of the text region detected by the localization model, depending on the experiment (see Section 4).
+GPT-4o
 GPT-4o [20] is OpenAI’s flagship language model in the Generative Pre-trained Transformer (GPT) family.
@@
 GPT models have demonstrated their versatility in information extraction and NER tasks, particularly in biomedical text analysis [17].
-Leveraging the superior performance of GPT in information extraction, we use GPT-4o to analyze the context of extracted texts from images in order to determine whether they contain PHI and which type of PHI is present.
+In our experiments, GPT-4o is configured to perform three distinct tasks: (1) text analysis, (2) text extraction, and (3) end-to-end image analysis.
+In the text analysis configuration, GPT-4o examines the context of extracted texts from images in order to determine whether they contain PHI and which type of PHI is present.
 The extracted texts from the text extraction module are provided to GPT-4o as a list.
 The system prompt consists of the definition of PHI and non-PHI texts and instructions to evaluate the aggregated context of the input text rather than analyzing each line in isolation.
-Additionally, we provide examples in the prompt to make the model aware of ambiguous cases, following the few-shot prompting technique [26].
-For instance, Age: or 65 contains a placeholder for age or hints at a number which could be age, but neither is considered PHI.
-To ensure consistent outputs, we constrain GPT-4o to produce a dictionary of mandatory fields through OpenAI’s pydantic tool, as outlined in Table 2.
+When GPT-4o acts as a text extractor, the prompts are structured to ensure that it returns the text visible in each cropped region predicted by YOLOv11.
+In the end-to-end image analysis scenario, GPT-4o is presented with a full image, prompted to perform text recognition, and tasked with analyzing each imprint identified in the image.
+It’s important to know that GPT-4o is not a dedicated object detection model, it does not return coordinates for the imprint.
+For the analysis task, we provide examples in the prompt to make the model aware of ambiguous cases, following the few-shot prompting technique [32].
+For instance, Age: or 65 refers to a placeholder for age or hints at a number which could be age, but neither is considered PHI.
+To ensure consistent outputs, we constrain GPT-4o to produce a dictionary containing mandatory fields through OpenAI’s pydantic tool, as outlined in Table 2.
 If the PHI is detected, the type field specifies one of the eight PHI classes or other if the PHI category is not included in the provided list.
@@
 Tag Data type Description
-type str One of [date, identifier, name, address, age, gender, phone_nr, email, other, non-phi]
+type str One of [date, identifier, patient_name, address, phone_nr, email, other, non-phi]
 raw_text str Original text extracted from the image
@@
 Specifically, we study the four setups illustrated in Figure 2.
-The first setup employs three distinct models, each dedicated to one of the modules described in Section 3.1.
-This serves as a baseline setup, where each model performs the task it is specifically designed for.
+The first setup employs three distinct models, each dedicated to one of the modules described in Section 3.
 In the second setup, we assess the robustness of EasyOCR’s built-in object detection capabilities by using it for both text localization and extraction tasks.
@@
 For each of the introduced setups, we run the pipeline five times on the RadPHI-test dataset to assess the non-deterministic behavior of GPT-4o in generating responses.
-The final output of the PHI pipeline is evaluated in two stages.
-In the first stage, the pipeline is treated as a binary classifier for detecting PHI.
-The objective is to predict whether or not the image contains any PHI.
-In the second stage, only images identified as containing PHI are considered.
-The goal here is to evaluate whether the PHI categories are correctly classified.
-This stage functions as a multi-label classification task across eight PHI classes, see Table 1.
-In both stages, we calculate the mean values of accuracy, precision, recall, and F1 score by aggregating results from five distinct runs for each setup.
+The final output of the PHI pipeline is evaluated at two levels: case and instance.
+At the case level, the focus is on whether the model can identify the presence of at least one PHI instance or a specific PHI class within the image.
+At the instance level, the evaluation checks whether the model accurately localizes every PHI instance or every instance of a specific class in the dataset.
+Performance is measured using recall, precision, false negatives (FN), and false positives (FP), which together provide insight into the model’s sensitivity in identifying PHI.
+A summary of the evaluation framework can be found in Table 3.
+In addition to the four pipeline variants, Microsoft’s open-source tool Presidio [16] is included in the benchmark.
+Presidio is designed for de-identifying images using NER models and regular expression patterns.
+We create two custom categories for patient names and identifiers using regular expressions, while also utilizing other built-in categories such as date, address, phone number, and email.
+Since Presdio is deterministic, we run it only once on our data and evaluate its performance.
+Our comparison with Presidio serves as a baseline for evaluating our LLM-based approaches.
 The setup that performs best on the RadPHI-test dataset is then further evaluated on the MIDI dataset using the same evaluation metrics.
+Level Evaluation type Interpretation Metric Dataset
+Case PHI presence Did the model correctly identify the presence of at least one PHI instance in the image? Recall, Precision, FN, FP RadPHI-test, MIDI
+Case PHI class presence Did the model correctly identify the presence of at least one PHI instance of class X in the image? Recall, Precision, FN, FP RadPHI-test, MIDI
+Instance PHI presence Did the model correctly localize every PHI instance in the dataset? Recall, Precision, FN, FP RadPHI-test(*), MIDI
+Instance PHI class presence Did the model correctly localize every PHI instance of class X in the dataset? Recall, Precision, FN, FP RadPHI-test(*), MIDI
+Table 3: Evaluation framework for PHI detection.
+The outputs are evaluated at two levels: Case-level, which considers whether the model detects the presence of at least one PHI instance or a specific PHI class in an image; and Instance-level, which evaluates the model’s ability to accurately localize each individual PHI instance or every instance of a specific PHI class across the dataset.
+(*) Instance-level evaluation is not performed in Setup 2 and Setup 4 because the coordinates are either not compatible to match with the ground-truth annotation (setup 2) or cannot be predicted (setup 4).
 5 Results
 5.1 Rad-PHI dataset
-(a) Precision
-(b) Recall
-(c) F1 Score
-(d) Accuracy
-Figure 7: Comparison of precision, recall, F1 score, and accuracy across four different setups, averaged from five separate runs of each setup obtained on the RadPHI-test dataset.
-Precision Recall F1 Acc
-Min Max Min Max Min Max Min Max
-Global PHI 0.9875 0.9995 0.9482 0.9972 0.9701 0.9983 0.9506 0.9972
-Date 0.9410 1.0000 0.8703 1.0000 0.9307 1.0000 0.9654 1.0000
-Patient Name 0.6913 0.9908 0.9025 0.9969 0.7827 0.9938 0.8786 0.9968
-Address 0.8725 1.0000 0.9571 0.9992 0.9128 0.9984 0.9563 0.9992
-Gender 0.9160 0.9967 0.6934 0.9483 0.8178 0.9714 0.9245 0.9864
-Email 0.9813 1.0000 0.8356 0.9964 0.9101 0.9985 0.9542 0.9990
-Phone Nr 0.9648 0.9985 0.8339 0.9977 0.9056 0.9981 0.9548 0.9990
-Age 0.9246 1.0000 0.6547 0.9913 0.7891 0.9957 0.9103 0.9978
-Identifier 0.8026 1.0000 0.8886 0.9797 0.8685 0.9897 0.9359 0.9952
-Table 3: Minimum and maximum scores of mean precision, mean recall, mean F1, and mean accuracy for each PHI category computed across all setups obtained on RadPHI-test.
+(a) Case-level precision
+(b) Case-level recall
+(c) Instance-level Precision
+(d) Instance-level Recall
+Figure 7: Comparison of precision and recall at case and instance level.
+Metric averages from five separate runs of each setup on the RadPHI-test dataset.
+Setups 2 and 4 are not evaluated at the instance level because of incompatible or absent predicted imprint coordinates.
+No. Images Mean Precision [FP] Mean Recall [FN]
+with PHI Worst Setup Best Setup Presidio Worst Setup Best Setup Presidio
+PHI Presence 779 0.9852 [11.4] 1.0000 [0] 0.9044 [54] 0.9017 [76.6] 0.9995 [0.4] 0.6560 [268]
+Date 267 0.9309 [17.8] 1.0000 [0] 0.5734 [61] 0.8689 [35] 0.9963 [1] 0.3071 [185]
+Patient Name 260 0.7532 [72.2] 1.0000 [0] 0.4102 [220] 0.8469 [39.8] 0.9923 [2] 0.5885 [107]
+Address 248 0.9256 [17.8] 0.9959 [1] 0.7634 [31] 0.8855 [28.4] 1.0000 [0] 0.4032 [148]
+Email 278 0.9950 [1.2] 1.0000 [0] 1.0000 [0] 0.8489 [42] 0.9849 [4.2] 0.2086 [220]
+Phone Nr 260 0.9843 [3.2] 1.0000 [0] 0.9250 [6] 0.7631 [61.6] 0.9908 [2.4] 0.2846 [186]
+Identifier 237 0.8301 [44.2] 1.0000 [0] 0.4065 [146] 0.8675 [31.4] 0.9949 [1.2] 0.4219 [137]
+Table 4: Worst and best performing setup based on mean precision and recall at case level computed across five different runs on RadPHI-test (1,000 images).
+The mean False-Positive (FP) and False-Negative (FN) values indicated in the brackets are computed for the corresponding setup across five different runs as well.
+Note that Presido operates deterministically, with the scores reported derived from a single run.
 The color-coding indicates the setup attaining the corresponding metric value, i.e.
 Setup 1, Setup 2, Setup 3, Setup 4.
+No. Instances Mean Precision [FP] Mean Recall [FN]
+with PHI Worst Setup Best Setup Presidio Worst Setup Best Setup Presidio
+PHI Presence 1550 0.9859 [21.6] 0.9869 [20.2] 0.3002 [1074] 0.9778 [34.4] 0.9808 [29.8] 0.3019 [1082]
+Date 267 0.9963 [1] 1.0000 [0] 0.2985 [188] 0.9655 [9.2] 0.9963 [1] 0.2809 [192]
+Patient Name 260 0.9642 [9.4] 0.9915 [2.2] 0.1511 [1117] 0.9731 [7] 0.9838 [4.2] 0.3923 [158]
+Address 248 1.0000 [0] 1.0000 [0] 0.7619 [70] 0.9798 [5] 1.0000 [0] 0.3750 [155]
+Email 278 1.0000 [0] 1.0000 [0] 0.9655 [2] 0.9496 [14] 0.9849 [4.2] 0.2014 [222]
+Phone Nr 260 0.9961 [1] 1.0000 [0] 0.8235 [15] 0.9823 [4.6] 0.9908 [2.4] 0.2692 [190]
+Identifier 237 0.9815 [4.4] 0.9913 [2] 0.2353 [312] 0.9671 [7.8] 0.9848 [3.6] 0.3038 [165]
+Table 5: Worst and best performing setup based on mean precision and recall at instance level computed across five different runs on RadPHI-test (1,000 images).
+The mean False-Positive (FP) and False-Negative (FN) values indicated in the brackets are computed for the corresponding setup across five different runs as well.
+Note that Presido operates deterministically, with the scores reported derived from a single run.
+The color-coding indicates the setup attaining the corresponding metric value, i.e.
+Setup 1, Setup 3.
 Setup Models Pipeline error rate Total time Prompt tokens Response tokens
-1 YOLOv11 + EasyOCR + GPT-4o $0.080.08$ % 3,64136413,6413 , 641 s 1,686,27716862771,686,2771 , 686 , 277 227,702227702227,702227 , 702
-2 EasyOCR + GPT-4o $0.100.10$ % 4,78547854,7854 , 785 s 1,704,06217040621,704,0621 , 704 , 062 259,819259819259,819259 , 819
-3 YOLOv11 + GPT-4o $0.280.28$ % 12,2951229512,29512 , 295 s 5,639,64956396495,639,6495 , 639 , 649 425,935425935425,935425 , 935
-4 GPT-4o $8.568.56$ % 3,43434343,$43$ , 434 s 3,249,45532494553,249,4553 , 249 , 455 250,577250577250,577250 , 577
-Table 4: Comparison of mean error rate, mean total time of analysis, the mean number of prompt and response tokens for 1,000 test cases across five runs for each setup evaluated on RadPHI-test.
+1 YOLOv11 + EasyOCR + GPT-4o $0$ % 4,14341434,1434 , 143 s 1,727,29017272901,727,2901 , 727 , 290 280,998280998280,998280 , 998
+2 EasyOCR + GPT-4o $0$ % 3,23532353,2353 , 235 s 1,744,12517441251,744,1251 , 744 , 125 298,087298087298,087298 , 087
+3 YOLOv11 + GPT-4o $1.61.6$ % 7,68576857,6857 , 685 s 6,152,36961523696,152,3696 , 152 , 369 474,746474746474,746474 , 746
+4 GPT-4o $5.185.18$ % 4,10441044,1044 , 104 s 3,530,30235303023,530,3023 , 530 , 302 287,229287229287,229287 , 229
+Baseline Presidio $0$ % $288288$ s N/A N/A
+Table 6: Statistics related to pipeline execution: mean pipeline error (failure) rate, mean total time of analysis, the mean number of prompt and response tokens for 1,000 test cases across five runs for each setup evaluated on RadPHI-test.
 The pipeline error rate is the percentage of images that fail to process due to server issues or the presence of banned words in the requests sent to GPT-4o.
-The total time is estimated as the sum of the processing and analysis time required for the entire dataset.
+The total time is estimated as the sum of the total processing and analysis time required for all 1,000 test cases.
 Figure 8: When text appears in multi-line blocks, out-of-the-box EasyOCR tries to fit bounding boxes that cover the entire text space.
 However, the bounding boxes might break the context consistency of the text.
-For example, Visit Date and 03-01-1989 are separated into two different crops, making it harder for the text analyzer to classify PHI content, particularly when a more fine-grain classification of dates is required, e.g., visit date or date of birth.
-The results of the four setups are summarized in Figure 7, showing the precision, recall, F1 score, and accuracy in classifying images for global PHI (i.e., whether the image has PHI or not) and the corresponding eight PHI categories.
-Table 3 compares the pipeline error rate, total time taken to process 1000 images, and the number of prompt and response tokens used by GPT-4o across four setups.
-The pipeline error rate indicates the percentage of images that couldn’t complete the process due to GPT-4o’s inability to handle the request or the presence of its banned words in the query.
-Overall, all four setups demonstrate strong performance in deciding if an image contains PHI imprints, with all evaluation metrics related to the Global PHI label exceeding $0.940.94$.
-Setup 2 (EasyOCR + GPT-4o) and setup 4 (GPT-4o) exhibit slightly lower performance compared to the other setups.
-However, their performance remains robust, with the overall accuracy at $0.96040.9604$ and $0.95060.9506$, respectively.
-Setup 3 (YOLOv11 + GPT-4o) achieves the highest performance, with all four metrics surpassing $0.990.99$.
-This indicates that the combination of YOLOv11 for text localization and GPT-4o for text extraction and analysis excels at identifying images containing PHI.
-The overall high scores in all evaluation metrics, even in lower-performing setups, suggest that detecting the presence of PHI containing imprints is for all pipeline setups a solvable task.
-Assessing the capability of different pipeline setups to perform fine-grained classification of PHI categories reveals strong performance deviations between the various setups.
-First, we observe that the PHI classes, such as date and address, are among the top-performing categories across all setups.
-This result is expected because both categories exhibit distinct semantic structures compared to other classes.
-On the other hand, gender information is the most challenging PHI class to detect, in particular when this information appears in abbreviated forms.
-Single-letter gender abbreviations often lack accompanying text like gender or sex, making detection more difficult.
-If the text extractor fails to recognize the exact character, the text analyzer is left without contextual clues to identify the PHI content and class.
-This issue is further complicated by potential confusion with single-character markers like L, R, A or P.
-Generally speaking, the bottleneck for fine-grained PHI category classification generally occurs during the text extraction step.
-When texts are incompletely extracted or when signaling texts are missed, it becomes significantly harder for the text analyzer to classify the content accurately.
-For example, Figure 8 illustrates how EasyOCR can break text continuity into separate bounding boxes, resulting in missing context and complicating the advanced analysis of PHI terms.
-Among the four setups, Setup 2 has one of the lowest performance levels.
-This outcome is expected, as the detection component of EasyOCR is not trained on medical image data, making it less robust compared to a dedicated YOLOv11 model.
-Even when EasyOCR is used solely for text extraction (Setup 1), the final pipeline performance lags slightly behind Setup 3 where GPT-4o is used for integrated text extraction and analysis.
-Upon manual examination of specific cases, EasyOCR demonstrates more errors in text extraction compared to using GPT-4o for this task.
-However, in a majority of cases, GPT-4o, when used as a text analyzer, is powerful enough to infer the intended meaning from improperly extracted text by EasyOCR.
-This capability results in a comparable performance between Setup 1 and 3.
-While Setup 3 offers a slightly better overall performance, it has a significant drawback: the total analysis time is more three times that of Setup 1.
+For example, Physician and Charlotte Lu are separated into two different crops, making it harder for the text analyzer to classify PHI content, particularly when a more fine-grain classification of names is required, e.g., physician or patient name.
+The results of the four setups are summarized in Figure 7, showing the precision and recall in classifying images for PHI and the corresponding six PHI categories at the case and instance level.
+The worst and best setups based on precision and recall at two levels are summarized in Table 4-5.
+PHI presence detection
+Setups 1 and 3 demonstrate across all categories and metrics strong performance in detecting PHI presence at both the case and instance level.
+Setup 2, which relies on localization and text extraction by EasyOCR, exhibits a varying performance for different PHI categories.
+Setup 4, which utilizes GPT-4o end-to-end, obtains the lowest precision and recall.
+Setups 1 and 3 consistently achieve the highest results, attaining a maximum mean precision of 1.0 ± 0.0 (Setup 1) and a mean recall of 0.9995 ± 0.0007 (Setup 3) at the case level.
+Notably, Setup 3, with a case-level mean recall of 0.9995 (FN: 0.4 ± 0.55), indicates that, on average, fewer than one image containing at least one PHI is missed.
+However, with stricter evaluation at the instance level, the best precision and recall values decrease to 0.9869 ± 0.0008 and 0.9808 ± 0.0012 for Setup 1, resulting in an average of 29.8 ± 1.93 instances containing PHI being missed in the best-performing setup.
+Setups 2 and 4 are only assessed at the case level and show inferior performance compared to Setups 1 and 3.
+Additionally, Presidio falls behind all setups in both case and instance evaluations.
+At the case level, Presidio achieves precision and recall rates of 0.9044 and 0.6560, respectively, which drop significantly to 0.3002 and 0.3019 when evaluated at the instance level.
+This indicates that the tool has difficulty completely identifying all instances of PHI within the images.
+PHI category detection
+The performance of Setup 1 and Setup 3 is notably superior across all categories, while Setup 2 and Setup 4 significantly lag behind.
+Furthermore, Setups 1 and 3 demonstrate consistent performance across all metrics, in contrast to Setups 2 and 4, which exhibit considerable variability.
+Specifically, in terms of precision, Setups 2 and 4 achieve perfect scores for email and phone numbers but struggle with patient names and identifiers.
+Conversely, regarding recall, email and phone numbers experience the worst performance in Setup 2 and 4.
+Emails and phone numbers have unique formatting and signaling words, which allow easier differentiation from other types and hence higher precision.
+However, when texts are incompletely extracted or when key signaling elements are missed, it becomes challenging for the text analyzer to classify the content accurately.
+In these instances, email addresses and phone numbers can be easily confused with patient names and identifiers, as the former often include names while the latter consist of numbers.
+One of the reasons for incomplete text extraction is illustrated in Figure 8.
+EasyOCR in Setup 2 can break text continuity into separate bounding boxes, resulting in missing context and complicating the subsequent analysis of imprints.
+Similar to Setup 2 and Setup 4, Presidio also shows a high precision for selected PHI categories (phone numbers, email addresses) but low metric values on all other categories.
+In terms of recall, Presidio does not yield satisfactory results across all categories.
+Tokens, latency, and pipeline error rate
+Table 6 compares the pipeline error rate, total time taken to process 1,000 images, and the number of prompt and response tokens used by GPT-4o across four setups.
+The pipeline error rate indicates the percentage of images that couldn’t complete the process due to GPT-4o’s inability to handle the request or the presence of blacklisted words in the query.
+The blacklisted words are often encountered by erroneous OCR outputs, especially noisy or misrecognized names, can inadvertently resemble words that trigger these safety mechanisms.
+However, these encounters are rare compared to the general request processing error.
+While Setup 3 offers a slightly better overall performance in a few PHI categories, it has a significant drawback: the total analysis time is more than twice that of Setup 1.
 This increase is primarily due to the latency associated with querying GPT-4o multiple times for extracting text from individual image crops.
@@
 These errors frequently arise from unstructured outputs or the generation of special tokens that can cause the model to fail to respond.
-While our experimental result underscores the impressive vision capabilities of GPT-4o, it is still inadequate to entirely replace specialized models for text localization and extraction tasks in a single pass.
-For further evaluation on the MIDI dataset, we choose Setup 1 over Setup 3 as the gain in performance of the latter is very limited whereas inference costs, latency and token costs are much more favorable for Setup 1.
+In comparison to all four setups, Presidio has the shortest processing time (on average $0.2880.288$s per case), which is approximately a factor 10 faster than the LLM-based pipeline from Setup 2 (on average $3.2353.235$s per case).
+In regard to inference speed, lightweight or rule-based models, which do not require significant computational resources, clearly outperform our GPT-4o-based setup.
+It is important to note that our PHI pipeline has not yet been optimized for throughput.
+Smaller models and optimized inference processes will allow us to significantly reduce computation time in the future.
 5.2 MIDI
-In comparison to our simulated RadPHI-test dataset, the MIDI dataset is more challenging due to its increased diversity and a larger number of imprints.
+For further evaluation on the MIDI dataset, we choose Setup 1 over Setup 3 as the gain in performance of the latter is very limited whereas inference costs, latency and token costs are much more favorable for Setup 1, see Table 6.
+In contrast to our simulated RadPHI-test dataset, the MIDI dataset offers a realistic benchmark which is more challenging due to its increased diversity and a larger number of imprints.
 Relying on signaling words might not be sufficient to detect PHI effectively.
 For example, an image series by itself may not be classified PHI, but if it includes a date element, as shown in Figure 6, it is considered PHI.
-Overall, the performance of Setup 1 on MIDI is comparable to that obtained on RadPHI-test, see Figure 9 (left).
-Notably, the global PHI detection has a recall rate close to 1.
-Key categories demonstrating strong performance include dates, patient names, and addresses.
-However, similar to findings from the RadPHI-test, the identification of gender remains particularly challenging, especially in the context of abbreviations.
-The precision of the identifier category is diminished due to misclassifications of numerous image- and study-related identifiers as PHI.
-These misclassifications arise from our initial prompt to GPT-4o, which states that all forms of identifiers are regarded as PHI.
-To address the needs of this scenario, we refine the prompt by explicitly stating that study or image-related identifiers should not be classified as PHI.
-The improvement in precision for the identifier category following this prompt adjustment is demonstrated in Figures 9 (right) and  10.
-This case serves as a compelling illustration of how LLMs can adapt to detect PHI based on specific criteria, rather than being confined to a fixed definition.
-Overall, this allows users to configure the PHI detection system more flexibly, for example, to meet study-specific criteria.
-(a) Same system prompt as in RadPHI-test
-(b) Modified system prompt for identifier
-Figure 9: Performance comparison of precision, recall, F1 score, and accuracy on the MIDI dataset using Setup 1: (a) performance with the RadPHI-test prompt and (b) performance with an adapted prompt for identifier classification.
+When employing the same prompt used for RadPHI-test, we obtain the case-level precision and recall of 0.9509 ± 0.0127 and 0.9986 ± 0.0030, respectively as shown in Table 7 (a).
+On average, less than one image (FN: 0.2 ± 0.45) with PHI is overlooked.
+At the instance level where precise matching is crucial, precision and recall decline to 0.8797 ± 0.0096 and 0.9639 ± 0.0047, respectively.
+Key categories demonstrating strong performance include phone numbers, patient names, and dates.
+The precision of the identifier category is low due to misclassifications of numerous image- and study-related identifiers as PHI.
+The lower precision observed in the identifier category can be attributed to the misclassification of many image- and study-related identifiers as PHI.
+This issue stems from our initial guidelines for RadPHI-test, which regard all types of identifiers as PHI.
+To correct this, we refine the prompt by explicitly stating that study or image-related identifiers should not be categorized as PHI.
+The resulting enhancement in precision for the identifier category is reflected in Table 7 (b) and Figure 9.
+The effects of the new prompt are evident not only in the improvement across all categories at the instance level but also in a minor reduction in performance at the case level for categories such as patient name and phone number.
+While this observation is intriguing, we do not delve deeply into further prompt optimization at this stage.
+However, it warrants investigation in our future work, where we plan to evaluate additional LLMs on a larger dataset and study this phenomenon.
+This prompt adaptation exemplifies how LLMs can be tailored to detect PHI in accordance with specific criteria, rather than adhering to a fixed definition.
+Ultimately, this flexibility allows users to configure the PHI detection system to better meet study-specific criteria.
+No. Images / (a) Same prompt as for RadPHI-test (b) Modified prompt
+Level Class Instances Precision [FP] Recall [FN] Precision [FP] Recall [FN]
+Case PHI 147 0.9509 [7.6] 0.9986 [0.2] 0.9697 [4.6] (*) 1.0000 [0.0]
+Date 116 0.9931 [0.8] 0.9914 [1.0] 1.0000 [0.0] (*) 0.9914 [1.0]
+Patient name 104 0.9962 [0.4] 0.9981 [0.2] 1.0000 [0.0] 0.9769 [2.4] (*)
+Address 36 0.8931 [4.0] 0.9222 [2.8] 0.9500 [1.8] (*) 0.9444 [2.0]
+Phone Nr 21 0.9193 [1.8] 0.9714 [0.6] 1.0000 [0.0] (*) 0.9143 [1.8] (*)
+Identifier 79 0.7427 [26.8] 0.9772 [1.8] 0.9398 [5.0] (*) 0.9848 [1.2]
+Instance PHI 549 0.8797 [72.4] 0.9639 [19.8] 0.9721 [15.4] (*) 0.9781 [12] (*)
+Date 207 0.9553 [8.8] 0.9082 [19] 0.9893 [2.2] (*) 0.9758 [5.0] (*)
+Patient name 161 0.9851 [2.4] 0.9863 [2.2] 0.9776 [3.6] 0.9752 [4.9]
+Address 39 0.8382 [7.0] 0.9282 [2.8] 0.9070 [3.8] (*) 0.9487 [2.0]
+Phone Nr 24 0.9073 [2.4] 0.9750 [0.6] 0.9648 [0.8] (*) 0.9083 [2.2] (*)
+Identifier 118 0.6589 [59.6] 0.9746 [3.0] 0.9325 [8.4] (*) 0.9814 [2.2]
+Table 7: Performance comparison on the MIDI dataset using Setup 1: (a) performance with the RadPHI-test prompt and (b) performance with an adapted prompt for identifier classification.
+Superscripts (*) indicate statistical significance (pairwise Mann–Whitney U test, p < 0.05)
 Figure 9: (Left) DICOM tag overlays with the DICOM viewer.
@@
 6 Discussion
-We investigate four different setups of an end-to-end pipeline for pixel-level PHI detection in medical images.
+Text localization
 For the localization step, we recommend using a dedicated object detection model instead of relying on the built-in localization component of a generic OCR model (Setup 2) or the localization capability of a multi-modal LLM (Setup 4).
-Our experimental results demonstrate that separating the localization step and using a localization model trained explicitly on the data reflecting the target domain ensures strong overall performance, even when out-of-the-box OCR tools, like EasyOCR, are used downstream for text extraction (Setup 1).
-End-to-end GPT-4o (Setup 4) has the drawback of not returning the bounding box coordinates of text regions, which is essential for subsequent PHI redaction.
+Our experimental results demonstrate that separating the localization step and using a localization model that can generalize well across different medical imaging modalities ensures strong overall performance, even when out-of-the-box OCR tools, like EasyOCR, are used downstream for text extraction (Setup 1).
+End-to-end GPT-4o (Setup 4) has the drawback of not returning the bounding box coordinates of text regions, which are essential for subsequent PHI redaction.
 Additionally, imprints are often incompletely identified and hallucinations occur much more frequently compared to other setups, resulting in lower performance.
-Moreover, there are more pipeline errors related to server issues or unstructured output, which prevents GPT-4o from generating a proper response.
-For the text extraction step, we observe that GPT-4o outperforms EasyOCR by having lower word-level and character-level errors.
-However, when evaluating the end-to-end pipeline performance, the improvement in PHI detection when switching to GPT-4o from EasyOCR (i.e, Setup 3 vs.
+Moreover, in this setup, more pipeline errors related to server issues or unstructured output occur, which prevents GPT-4o from generating proper responses.
+Text extraction
+When evaluating the OCR performance separately, GPT-4o outperforms EasyOCR by having lower word-level and character-level errors.
+However, for the end-to-end pipeline performance, the improvement in PHI detection when switching to GPT-4o from EasyOCR (i.e, Setup 3 vs.
 Setup 1) is minimal.
-This is because minor OCR errors produced by EasyOCR can easily be circumvented by the robust language model in the text analysis step.
+This is because minor OCR errors produced by EasyOCR can easily be compensated by the robust language model in the text analysis step.
 While GPT-4o offers superior performance in text extraction, the overall gain in pipeline performance is modest and accompanied by higher latency and increased costs associated with token generation.
 Therefore, EasyOCR is preferred in our study.
-For text analysis, GPT-4o is the default and only choice across four different setups.
+Text analysis
+For text analysis, GPT-4o is the default and only choice across all four different setups.
 The model excels at understanding text content and is particularly efficient in identifying PHI and further classifying it into sub-categories.
 Additionally, it can accurately discern context even when there are slight OCR errors in the extracted text.
-For instance, Figure 11 illustrates OCR spelling errors related to dates, identifiers, and gender.
+For instance, Figure 10 illustrates OCR errors related to dates, identifiers, and names.
 The model is capable of implicitly correcting these errors and reasoning based on the corrected text.
-Furthermore, one of the strengths of LLM is its flexibility, which allows for adjustments in instructions through prompts and configurations of output structures.
-This adaptability makes it easy to redefine PHI and associated tasks across various studies.
+Furthermore, an outstanding strength of LLM-based text analysis is the possibility to easily configure rules, sequential steps, and the output structure of the analysis through prompts.
+This adaptability makes it easy to redefine PHI and associated tasks across various studies or to accommodate different guidelines.
 Figure 10: (Left) An image from the RadPHI-test dataset with simulated imprints.
@@
 Despite the presence of spelling errors in the extracted text, GPT-4o effectively recognizes and interprets the content.
+Generalization
 A key concern for any AI-based system is its ability to generalize to unseen data.
-In light of this, we explore the use of pre-trained or foundational models with minimal fine-tuning, applicable solely to the localization model.
-Evaluation on the hold-out MIDI dataset has demonstrated that the YOLO model, despite being fine-tuned on a different dataset (RadPHI-train), can perform effectively on new datasets containing modalities not encountered during training, including PET, mammography, and ultrasound.
-Furthermore, the LLM demonstrates adaptability to diverse contexts, as evidenced by our results on the MIDI dataset with modified prompting.
-This underscores a valuable use case in which users can interactively adjust the prompt to redact or retain specific information in the images, all without the need to develop a new classification model.
-Overall, these findings emphasize the potential of leveraging existing models to enhance performance across various medical imaging domains.
-In terms of limitations, the expansion of GPT-4o’s capabilities beyond basic text analysis to the interpretation of small image crops and, ultimately, entire images reveals interesting findings.
-While the model demonstrates exceptional performance when analyzing image crops (Setup 3), it encounters a significant increase in technical errors and hallucinations when processing complete images (Table 3).
+In context of PHI detection, we believe the main source of variability stems from the text representation such as font size, color, and background, as well as the content itself, which may include PHI in various languages or formats not explicitly defined within the default PHI criteria.
+We address this by leveraging pretrained models whenever possible, requiring minimal fine-tuning.
+By choosing pretrained models, we preserve the models’ learned abstractions and only guide their reasoning towards specific criteria.
+For example, in our optimal setup (Setup 1), the only model that requires fine-tuning is the YOLOv11 model for imprint localization.
+By covering key imaging modalities, generating diverse imprints and utilizing numerous data augmentations methods during training, we achieve during inference that the YOLOv11 model performs robustly in unseen modalities that are not encountered during training, including PET, mammography, and ultrasound as part of the MIDI dataset.
+The other two models, namely EasyOCR and GPT-4o, work robustly with either no fine-tuning or only minor adjustments in the case of GPT-4o.
+On the MIDI dataset, which we have little control over the format of the imprints that appear in the images exported from the viewer.
+The results show that even without optimizing the prompt to tailor the formatting or representation, we achieve a decent recall rate of 0.9639, missing on average 19.8 out of 549 PHI instances.
+The precision is low, 0.8797, with 72.4 false positives on average due to unadjusted guidelines for classification.
+However, after adjusting the prompt to define specific criteria for inclusion and exclusion, we reduce the false positives to 15.4, improving the precision to 0.9721.
+This adaptation occurs only in the text analysis to conform to the new guidelines we select, while the text localization and extraction processes continue to perform well, even with the new dataset.
+While it is true that the LLM may struggle with new guidelines not covered in the initial prompt, modifying the PHI classification criteria is relatively straightforward using natural language.
+This stands in contrast with the complex engineering needed to customize rule-based classification systems or to fine-tune models with fixed rules.
+Additionally, it is less complicated to extend PHI classification to encompass diverse languages with LLMs, which is an advantage of using LLMs for this purpose.
+Capabilities of multi-modal LLM
+The expansion of GPT-4o’s capabilities beyond basic text analysis to the interpretation of small image crops and, ultimately, entire images reveals interesting findings.
+While the model demonstrates exceptional performance when analyzing image crops (Setup 3), it encounters a significant increase in technical errors and hallucinations when processing complete images in Setup 4 (Table 6).
 This inefficiency results in incomplete and inaccurate detection of PHI.
@@
 In our future work, we plan to investigate and benchmark various prompting strategies and assess the model’s robustness in languages other than English.
-Additionally, given the high level of privacy required for medical data, it is essential to either fine-tune a LLM or host it in-house for application development.
-Smaller LLMs may be particularly interesting to examine due to their ease of deployment and the ability to customize them for specific tasks.
-Furthermore, we plan to benchmark our pipeline against commercially available products [8, 7, 9, 10] currently available in the market.
+Additionally, given the high level of privacy required for medical data, it is essential to either fine-tune a LLM or host it in-house.
+Smaller LLM architectures are here of particular interest due to their ease of deployment, increased inference speed, reduced costs and the ability to customize them for specific tasks.
+Limitations and future work
+In our study, we focus on identifying the most effective setup for constructing a PHI pipeline, without delving deeply into benchmarking specific model types or variants to determine which offers the best performance.
+By optimizing the models used in each module separately, we acknowledge that our findings may not necessarily apply to particular models.
+For instance, we believe that some commercial OCR tools can process images effectively without requiring further fine-tuning, suggesting that Setup 2 may perform comparably to Setup 1.
+Moreover, our evaluation does not include testing our proposed approaches on real-world data containing PHI.
+In general, it is challenging to obtain medical image data containing actual PHI imprints for research and analysis in a privacy compliant way.
+Regulatory guidelines require that detected PHI is removed or de-identified by the data processor.
+In view of this, the MIDI dataset is created by utilizing an industry-standard medical image viewer to overlay synthetically generated PHI on medical images.
+Exporting image data directly from this viewer allows us to generate test data that is very close to reality.
+We recognize that real-world datasets are very heterogeneous and may comprise a diverse range of modalities and scanners.
+Hence, we plan to further extend our benchmark in future work to incorporate a wider variety of model types and curate additional datasets for validation, and extend the language beyond English.
 7 Conclusion
@@
 Appendix A Results on RadPHI-test
-Class Precision Recall F1 Accuracy
-Global PHI 0.9988 ± 0.0000 0.9904 ± 0.0016 0.9946 ± 0.0008 0.9908 ± 0.0013
-Date 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000
-Patient Name 0.9493 ± 0.0030 0.9961 ± 0.0001 0.9722 ± 0.0016 0.9852 ± 0.0008
-Address 0.9976 ± 0.0022 0.9992 ± 0.0018 0.9984 ± 0.0017 0.9992 ± 0.0008
-Gender 0.9967 ± 0.0030 0.7414 ± 0.0062 0.8512 ± 0.0059 0.9363 ± 0.0015
-Email 1.0000 ± 0.0000 0.9509 ± 0.0033 0.9749 ± 0.0017 0.9864 ± 0.0009
-Phone Nr 0.9923 ± 0.0000 0.9923 ± 0.0000 0.9923 ± 0.0000 0.9960 ± 0.0000
-Age 1.0000 ± 0.0000 0.9266 ± 0.0051 0.9619 ± 0.0028 0.9812 ± 0.0013
-Identifier 0.9948 ± 0.0019 0.9747 ± 0.0030 0.9847 ± 0.0018 0.9928 ± 0.0008
+Class Case Precision Case Recall Instance Precision Instance Recall
+PHI Presence 1.0000 ± 0.0000 0.9923 ± 0.0000 0.9869 ± 0.0008 0.9808 ± 0.0012
+Date 1.0000 ± 0.0000 0.9963 ± 0.0000 0.9963 ± 0.0000 0.9963 ± 0.0000
+Patient Name 0.9778 ± 0.0032 0.9815 ± 0.0017 0.9642 ± 0.0033 0.9731 ± 0.0000
+Address 0.9952 ± 0.0018 1.0000 ± 0.0000 0.9857 ± 0.0045 1.0000 ± 0.0000
+Email 1.0000 ± 0.0000 0.9496 ± 0.0000 1.0000 ± 0.0000 0.9496 ± 0.0000
+Phone Nr 1.0000 ± 0.0000 0.9908 ± 0.0021 1.0000 ± 0.0000 0.9908 ± 0.0021
+Identifier 0.9916 ± 0.0000 0.9949 ± 0.0035 0.9815 ± 0.0023 0.9848 ± 0.0023
 Table A1: PHI detection performance of setup 1 (YOLOv11 + EasyOCR + GPT-4o) on RadPHI-test.
-Class Precision Recall F1 Accuracy
-Global PHI 0.9875 ± 0.0011 0.9656 ± 0.0013 0.9764 ± 0.0009 0.9604 ± 0.0016
-Date 1.0000 ± 0.0000 0.8703 ± 0.0063 0.9307 ± 0.0037 0.9654 ± 0.0017
-Patient Name 0.8480 ± 0.0039 0.9308 ± 0.0027 0.8874 ± 0.0023 0.9385 ± 0.0014
-Address 0.9812 ± 0.0022 0.9694 ± 0.0036 0.9752 ± 0.0023 0.9878 ± 0.0011
-Gender 0.9965 ± 0.0032 0.6934 ± 0.0034 0.8178 ± 0.0029 0.9245 ± 0.0012
-Email 0.9991 ± 0.0019 0.8356 ± 0.0042 0.9101 ± 0.0031 0.9542 ± 0.0015
-Phone Nr 0.9909 ± 0.0032 0.8339 ± 0.0032 0.9056 ± 0.0027 0.9548 ± 0.0013
-Age 0.9929 ± 0.0026 0.6547 ± 0.0045 0.7891 ± 0.0032 0.9103 ± 0.0010
-Identifier 0.8492 ± 0.0061 0.8886 ± 0.0056 0.8685 ± 0.0038 0.9361 ± 0.0019
+Class Case Precision Case Recall
+PHI Presence 0.9852 ± 0.0007 0.9743 ± 0.0016
+Date 0.9983 ± 0.0023 0.8689 ± 0.0065
+Patient Name 0.8547 ± 0.0011 0.9323 ± 0.0058
+Address 0.9942 ± 0.0022 0.9750 ± 0.0018
+Email 1.0000 ± 0.0000 0.8489 ± 0.0025
+Phone Nr 0.9908 ± 0.0000 0.8254 ± 0.0021
+Identifier 0.8301 ± 0.0027 0.9114 ± 0.0000
 Table A2: PHI detection performance of setup 2 (EasyOCR + GPT-4o) on RadPHI-test.
-Class Precision Recall F1 Accuracy
-Global PHI 0.9995 ± 0.0007 0.9972 ± 0.0006 0.9983 ± 0.0005 0.9972 ± 0.0008
-Date 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000
-Patient Name 0.9908 ± 0.0021 0.9969 ± 0.0017 0.9938 ± 0.0008 0.9968 ± 0.0004
-Address 1.0000 ± 0.0000 0.9959 ± 0.0001 0.9980 ± 0.0000 0.9990 ± 0.0000
-Gender 0.9957 ± 0.0000 0.9483 ± 0.0063 0.9714 ± 0.0033 0.9864 ± 0.0016
-Email 1.0000 ± 0.0000 0.9964 ± 0.0000 0.9985 ± 0.0007 0.9990 ± 0.0000
-Phone Nr 0.9985 ± 0.0021 0.9977 ± 0.0021 0.9981 ± 0.0014 0.9990 ± 0.0007
-Age 1.0000 ± 0.0000 0.9913 ± 0.0018 0.9957 ± 0.0009 0.9978 ± 0.0004
-Identifier 1.0000 ± 0.0000 0.9797 ± 0.0082 0.9897 ± 0.0042 0.9952 ± 0.0019
+Class Case Precision Case Recall Instance Precision Instance Recall
+PHI Presence 0.9954 ± 0.0007 0.9995 ± 0.0007 0.9859 ± 0.0006 0.9778 ± 0.0021
+Date 1.0000 ± 0.0000 0.9655 ± 0.0041 1.0000 ± 0.0000 0.9655 ± 0.0041
+Patient Name 1.0000 ± 0.0000 0.9923 ± 0.0047 0.9915 ± 0.0017 0.9838 ± 0.0032
+Address 0.9959 ± 0.0029 0.9798 ± 0.0040 0.9959 ± 0.0029 0.9798 ± 0.0040
+Email 1.0000 ± 0.0000 0.9849 ± 0.0053 1.0000 ± 0.0000 0.9849 ± 0.0053
+Phone Nr 0.9961 ± 0.0000 0.9823 ± 0.0034 0.9961 ± 0.0000 0.9823 ± 0.0034
+Identifier 1.0000 ± 0.0000 0.9755 ± 0.0046 0.9913 ± 0.0000 0.9671 ± 0.0046
 Table A3: PHI detection performance of setup 3 (YOLOv11 + GPT-4o) on RadPHI-test.
-Class Precision Recall F1 Accuracy
-Global PHI 0.9929 ± 0.0015 0.9482 ± 0.0030 0.9701 ± 0.0013 0.9506 ± 0.0021
-Date 0.9410 ± 0.0072 0.9513 ± 0.0050 0.9461 ± 0.0030 0.9722 ± 0.0017
-Patient Name 0.6913 ± 0.0147 0.9025 ± 0.0146 0.7827 ± 0.0082 0.8786 ± 0.0063
-Address 0.8725 ± 0.0048 0.9571 ± 0.0083 0.9128 ± 0.0039 0.9563 ± 0.0021
-Gender 0.9160 ± 0.0090 0.7560 ± 0.0095 0.8284 ± 0.0081 0.9278 ± 0.0024
-Email 0.9813 ± 0.0019 0.9034 ± 0.0075 0.9407 ± 0.0038 0.9696 ± 0.0017
-Phone Nr 0.9648 ± 0.0063 0.8697 ± 0.0117 0.9148 ± 0.0068 0.9597 ± 0.0031
-Age 0.9246 ± 0.0089 0.9431 ± 0.0067 0.9337 ± 0.0038 0.9666 ± 0.0022
-Identifier 0.8026 ± 0.0109 0.9550 ± 0.0073 0.8722 ± 0.0084 0.9359 ± 0.0050
+Class Case Precision Case Recall
+PHI Presence 0.9918 ± 0.0023 0.9017 ± 0.0066
+Date 0.9309 ± 0.0087 0.8966 ± 0.0146
+Patient Name 0.7532 ± 0.0084 0.8469 ± 0.0202
+Address 0.9256 ± 0.0184 0.8855 ± 0.0175
+Email 0.9950 ± 0.0054 0.8633 ± 0.0171
+Phone Nr 0.9843 ± 0.0090 0.7631 ± 0.0169
+Identifier 0.8809 ± 0.0230 0.8675 ± 0.0238
 Table A4: PHI detection performance of setup 4 (GPT-4o) on RadPHI-test.
 Appendix B Results on MIDI
-Class Precision Recall F1 Accuracy
-Global PHI 0.9900 ± 0.0034 0.9962 ± 0.0035 0.9931 ± 0.0014 0.9890 ± 0.0022
-Date 1.0000 ± 0.0000 0.9948 ± 0.0077 0.9974 ± 0.0039 0.9970 ± 0.0045
-Patient Name 0.9735 ± 0.0076 0.9885 ± 0.0105 0.9809 ± 0.0059 0.9800 ± 0.0061
-Address 0.9631 ± 0.0291 0.9889 ± 0.0152 0.9755 ± 0.0110 0.9910 ± 0.0042
-Gender 0.9965 ± 0.0079 0.9492 ± 0.0000 0.9722 ± 0.0038 0.9840 ± 0.0022
-Phone Nr 0.9723 ± 0.0253 0.9619 ± 0.0399 0.9663 ± 0.0138 0.9930 ± 0.0027
-Age 0.9637 ± 0.0133 0.9606 ± 0.0083 0.9621 ± 0.0075 0.9750 ± 0.0050
-Identifier 0.7622 ± 0.0093 0.9899 ± 0.0106 0.8612 ± 0.0088 0.8740 ± 0.0082
+Class Case Precision Case Recall Instance Precision Instance Recall
+PHI Presence 0.9509 ± 0.0127 0.9986 ± 0.0030 0.8797 ± 0.0096 0.9639 ± 0.0047
+Date 0.9931 ± 0.0039 0.9914 ± 0.0000 0.9553 ± 0.0081 0.9082 ± 0.0048
+Patient Name 0.9962 ± 0.0052 0.9981 ± 0.0043 0.9851 ± 0.0055 0.9863 ± 0.0081
+Address 0.8931 ± 0.0305 0.9222 ± 0.0232 0.9851 ± 0.0055 0.9863 ± 0.0081
+Phone Nr 0.9193 ± 0.0186 0.9714 ± 0.0261 0.9073 ± 0.0189 0.9750 ± 0.0228
+Identifier 0.7427 ± 0.0188 0.9772 ± 0.0057 0.6589 ± 0.0167 0.9746 ± 0.0060
 Table B1: PHI detection performance of setup 1 (YOLOv11 + EasyOCR + GPT-4o) on MIDI using the same prompt as RadPHI-test
-Class Precision Recall F1 Accuracy
-Global PHI 0.9987 ± 0.0028 0.9924 ± 0.0029 0.9955 ± 0.0017 0.9930 ± 0.0027
-Date 0.9983 ± 0.0038 0.9931 ± 0.0038 0.9957 ± 0.0030 0.9950 ± 0.0035
-Patient Name 0.9923 ± 0.0124 0.9808 ± 0.0000 0.9865 ± 0.0062 0.9860 ± 0.0065
-Address 0.9472 ± 0.0255 0.9889 ± 0.0152 0.9675 ± 0.0181 0.9880 ± 0.0067
-Gender 0.9965 ± 0.0078 0.9627 ± 0.0142 0.9793 ± 0.0099 0.9880 ± 0.0057
-Phone Nr 0.9735 ± 0.0391 1.0000 ± 0.0000 0.9862 ± 0.0204 0.9970 ± 0.0045
-Age 0.9639 ± 0.0079 0.9697 ± 0.0000 0.9668 ± 0.0040 0.9780 ± 0.0027
-Identifier 0.9920 ± 0.0118 0.9342 ± 0.0165 0.9621 ± 0.0087 0.9709 ± 0.0066
+Class Case Precision Case Recall Instance Precision Instance Recall
+PHI Presence 0.9697 ± 0.0035 1.0000 ± 0.0000 0.9721 ± 0.0032 0.9781 ± 0.0043
+Date 1.0000 ± 0.0000 0.9914 ± 0.0000 0.9893 ± 0.0064 0.9758 ± 0.0076
+Patient Name 1.0000 ± 0.0000 0.9769 ± 0.0129 0.9776 ± 0.0035 0.9752 ± 0.0116
+Address 0.9500 ± 0.0116 0.9444 ± 0.0278 0.9070 ± 0.0092 0.9487 ± 0.0256
+Phone Nr 1.0000 ± 0.0000 0.9143 ± 0.0213 0.9648 ± 0.0197 0.9083 ± 0.0186
+Identifier 0.9398 ± 0.0136 0.9848 ± 0.0057 0.9325 ± 0.0115 0.9814 ± 0.0038
 Table B2: PHI detection performance of setup 1 (YOLOv11 + EasyOCR + GPT-4o) on MIDI using the modified prompt with focus on eliminating false positives in the identifier category.
@@
 Appendix D Examples of pipeline output
-Figure D1: Comparison of outputs from four setups: Setups 1, 3, and 4 demonstrate similar response patterns, with Setup 3 accurately identifying the gender.
-In contrast, Setup 2 exhibits incomplete text extraction, missing the patient’s name and gender.
-Figure D2: Comparison of outputs from four setups: Setup 1 and Setup 3 accurately classify PHI categories despite overlapping address image regions.
-Setup 2 fails to recognize the age class, while Setup 4 illustrates model hallucination, generating responses that are not based on the provided input.
+Figure D1: Comparison of outputs from four setups: Setups 1, 3, and 4 demonstrate similar response patterns. In contrast, Setup 2 exhibits duplicate text extraction, resulting in similar entries.
+Figure D2: Comparison of outputs from four setups: Setup 1, 2, and 3 accurately classify PHI category while Setup 4 illustrates model hallucination, generating responses that are not based on the provided input.
 Figure D3: Comparison of responses from the same case across five distinct runs.